A Bibliography of Publications in Notices of the American Mathematical Society (2000–2009)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

15 August 2024 Version 1.00

Title word cross-reference

(τ) [Lub05]. $45.00 [Sha05]. $49.00 [Sha05]. abc [GT02]. E_{8} [Vog07]. G_{2} [Agr08]. ∞ [Lur08]. L [Boo08]. p [CK05]. SL_{2}(R) [Kis07a]. U(N) [KTT09].

-adic [CK05]. -category [Lur08]. -function [Boo08].

1 [Sha05]. 1-56881-136-5 [Sha05]. 1-56881-211-6 [Sha05]. 100th [Jac04d]. 120-cell [Sti01].

2000 [Ano00a, Ano00b, Ano00c, Ano01a]. 2001 [Ano01b, Ano01d, Ano01c, Ano01e, Ano01f].

2002 [Ano02a, Ano02b, Ano02d, Ano02c, Ano02e, Ano03, FRS03, Jac03d]. 2003 [Ada05]. 2004 [Ano04f, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Jac06e]. 2005 [Jac05e]. 2006 [Ano06c, Ano06a, Ano06d, Ano06b, Ano07i, Jac06c, Jac06a, VBKR07]. 2007 [Ano07a, Ano07b, Ano07c, Ano07d, Ano07f, Ano07e, Ano07g, Ano07m, Ano07h]. 2008 [Dys09]. 2009 [Ano08a, Ano08c, Ano09a, Ano09c, Ano09b, Ano09i, Ano09j, Ano09d, Ano09f, Ano09e, Ano09g].

3-manifolds [And04, Mil03].

45th [Ano01g].

5 [Sha05].

88i [Row01].

98a [Dav01].

A. [Dav01]. Abel [Ano04f, Ano07m, Ano09j, Ano05c, Art05, Jac03e, RS07]. academic [Leo08]. according [Har00]. Action [Pet10]. Addition [AO01]. address [Art05]. adic [CK05]. Adleman [Ano03]. Advanced [Lan00a]. Adventures [Leo08]. age [Tan06]. aid [Fow05]. Akaike [Ano06c]. Alan [Coo06]. Alexander [Sch08]. Alexandre [Jac04a, Jac04b]. Alexandrovna [FLM^{+}04]. Alfred [Wel07]. Algebra [KN08, Car05, Cox05, Guo09, Zel07, Ano09b, Art09]. algorithm [Cro09]. algorithms [Jor03, McG01]. Alhambra [Grü06]. American [Bat08, Bat09, Jac05a]. Americans [Goe06]. AMS [Ano00a, Ano09a, Art05, Jac09a, Jac09b, Lan05, Zit09]. analogy [Kri05]. analysis [Art00, CHW09, HK09, IK00, Kad08, McG01, Sem03a, Sem03b, Tao01, Tre00]. André [KS09, Kri05]. Andrew [Ano09h]. Andrews [Ano09k]. Anneli [Sau00]. anniversary [Ale07, Jac04g]. annotated [Hod01]. answered [Knu02]. appelé [Jac04a, Jac04b]. Applications [Per00, Mar06]. applied [Bar01, Cox05]. approach [Moo09]. Archimedes [Gra08]. archives [BCLP03]. Arild [Per05]. arithmetic [AO01]. Armand [ABC^{+}04]. Arnold [Ano08a, Jac01a, JS03]. Arsen’evna [JLM03]. Artful [dSLDS03]. Arthur [Jac05d]. ascendance [Bat09]. Asian [Goe06]. asks [Tan06]. Astrophysics [KN08]. Asymptotic [Sha01]. Atiyah [Ano04f, JW06, RS05]. Atlas [Har04]. Atle [Ano09l]. atomic [PC05]. attractor [Rue06]. attractors [Coe09]. Aumann [Ney06]. authors [Hod01]. autodidact [Fow05]. award [Jac04i, Ano02a, Ano03, Jac00e]. awarded [Ano06a, Ano06b, Ano07l, Ney06]. AWM [Ano01g]. axioms [Myc06].

B [CJW07]. Bôcher [Ano02b, Bat09]. Babylonian [Fri08]. back [DS01, Row01]. bad [Eps06]. Bailey [Sha05]. ball [DM08]. Barcelona [Jac00b]. Bartle [KS04]. basis [Mar06, Stu05]. Bass [Jac01c, Jac03c]. Batchelor [Bar01]. Baxter [Guo09]. be [Gol09b]. beauty [Ada05, Ada05]. bedroom [Aus09]. been [Cas07]. Beilinson [Ano00d]. Bers [Jac07a]. best [Sus09]. bestseller [Row01]. Better [BJK06]. between [Kob07, Tao01]. billiard [Sin04]. billiards [PL06]. biographer [Daw06]. biological [Gro00]. biology [Kar02, MS05, Ree04]. Birds [Dys09]. Birkhoff [Ano09a]. Birman [Ano07k]. birthday [Jac03a]. Bit [Cas06a]. Black [Bra02, Was08]. blind [Jac02c]. block [Ber02]. blow [CF03]. blow-up [CF03]. Blueprint [Cas06a]. body [Mon01]. Bohr [PC05]. Book [Coo06, Dav01, Hod06, Mor05a, Sha05, Art09, Aus09, BDK^{+}09, Bla09, Coe09, Gun09, Har04, Joh09, Moo09, Sas09, Sha09, Sus09, Gol09b]. Books [Har04]. Boor [Ano05a]. Borel [ABC^{+}04]. Boris [Lub04]. Borwein [Sha05]. Bott [Jac01b, Tu06]. boundary [Fri00]. bowl [CMP04]. brain [FS05, Gro00]. brane [Moo05]. Breaking [Grä09a]. breakthrough [Bor03]. Brownian [Sla02]. Bulgaria [Sau03]. bundle [BGPG07].

C [KGG^{+}04, dSLDS03]. Cahn [Sau00]. cake [BJK06]. calculus [Bla09]. Call [Ano01g]. Cambridge [Dav01]. Cameron [Abh02]. Cannonballs [Hal00]. Cantor [Dev04b]. cardinal [Ste07]. Carleson [RS07]. Carlitz [Hay01]. Carlo [BBS^{+}00]. Cartan [Aud09, Jac04d]. case [Sas09]. Casimir [KTT09]. Castelnuovo [BG09]. Casti [Dav01]. category [Lur08]. celebrates [Ale07, Jac03a, Jac04g]. celebration [Art05, Gra08]. cell [Sti01]. Cells [Gun06]. centennial [Gra08]. century [Grä07, Kad08, Lan00a]. challenge [Jaf06]. challenges [Kar02, Neu01]. Champaign [Gra03]. championship [CMP04]. change [Gol09a]. Changing [Dur09]. chaos [Lyu00, Rud08]. character [Vog07]. Charles [Jac07a]. checklist [Hod01]. Chern [Ano04g, Jac05b]. choose [Gel09]. chooses [Gel09]. chromatic [HM07]. Circle [Ste03]. circles [Tan06]. classical [Mez07]. classification [Asc04, Mil03]. Claude [GBC^{+}02]. Clerk [Sil08]. Climate [Gol09a]. cluster [Zel07]. coarse [Roe06]. codes [Ber02, Elk00a, Elk00b]. Cohomology [Car05]. Cole [Ano00b, Ano02c, Ano09b, BCF^{+}00]. collapse [CF03]. collection [Fri08]. color [Gon08]. Combinatorial [Dev04a, Dei00]. combinatorics [Ios01, Tao01]. come [DS01]. coming [Tan06]. Comme [Jac04a, Jac04b]. common [Sas09]. communication [Min05]. Communications [Lan00a]. Communities [POM09]. commutative [Car05]. commutator [Cal08]. compact [Mez07]. compacton [Ros05]. companion [BDK^{+}09]. Company [Har04]. Competitions [Ano01g]. complex [Dev04b, Gin09, HK09, IK00]. complexity [Ada05]. Computable [Mil08b]. computation [EE03]. computations [Car05]. Computer [Coo06]. Computing [Blu04, BC06]. Conant [Ano01b, Ano02d, Ano04a, Ano07a, Ano09c]. concept [HK09]. concise [Row01]. Conference [Ano01g]. confusion [WAD09]. Congressman [Ano07j]. conjecture [Mil03, Szp03, Mor05a]. Conjectures [Jac06b]. Connected [Zit09]. connections [Tao01]. Consensus [Jac06b]. Contest [Ano01g]. continuum [Woo01a, Woo01b, Woo02]. contributions [JK07]. control [PC05]. Conversations [Kir02]. converse [DGPV07]. Cooley [MR01]. Correction [Woo02]. Corrections [Ano01g]. Cotlar [Hor09]. could [Cho06]. counting [AO01]. Coverand [Cas06a]. Coxeter [EGMW03, Gun06]. creating [Sun08]. creative [Sas09]. Crighton [Bar01]. cross [Lab08]. crunchers [Gol09b]. cryptography [Kob07]. Cryptologia [Ano01g]. crystal [DM08]. Crystals [Sun08]. Csicsery [Cas08]. curse [KS05]. Curvature [Ios01]. curve [Don05, Mik07]. curvelet [Can03]. curves [Vak03]. cusp [Wri04]. cut [BJK06].

D [BCF^{+}00, Har04, RS06, Spr00]. D. [Row01]. daemon [Wel07]. Dantzig [CJW07]. data [Gol09b, PS03, Lan00b]. Davey [DLM^{+}07]. David [Sha05, Ano05d, Bar01, Jac03b]. Davis [Jac08b]. Dawson [Dav01]. dealing [PS03]. Deane [Fin05]. decompositions [KTT09]. Dehn [Daw02]. density [Mor05b]. Departments [Ano01g]. DePauli [Dav01]. depth [PFTY09]. derivative [OT09]. Descartes [Ser08]. destiny [Sus09]. Deuring [Sto06]. diagrams [Mar06, RSW06]. different [Cul02, Jac05a]. Differential [Aud08, Mar06]. difficulties [Cas04]. diffusions [SC01]. dilemmas [Dav01]. dimension [Lip09]. dimensionality [KS05]. dimensions [Cas04]. dimer [KO05]. director [Jac09a]. Dirk [DTRB01]. discovery [RMM08]. Discretizing [HS04]. Disks [CM03]. distance [Abh02]. distances [IK00]. Distinguished [Ano02a]. diversions [Aus09]. Dixmier [Sch09]. do [Cul02, Gel09, Hod01, Kri04]. dodecahedral [Wee04]. domain [CG08]. Donald [Jac09a, KGG^{+}04]. Donaldson [Ano08b, Ano09i, Jac06c]. doors [Jac00c]. Doting [Gol09b]. Douady [Sch09]. Double [Wri04, CM03]. Dover [Row01]. down [Aus05]. dual [Gin09]. dummies [Sch09]. Dynamical [Rue02, Cou06, Kop00]. dynamics [BMZ05, Dev04b, Khe05]. early [Rie02]. easy [GT02]. ecology [Neu01]. Editor [DLM^{+}07]. education [McC07]. Edward [Jac00a]. Eells [Tol08]. Efron [Jac07b]. Einstein [Dys09, DS01, JW06]. Eisenbud [Ano05d, Jac03b]. electronic [Jac00f]. Elements [Gra08]. elephant [Bis04]. elliptic [Och09]. Elwood [GBC^{+}02]. emerging [Erm04, Tao01]. Emma [Bri07]. Emmy [Art09]. Encounter [Ber00a, Ber00b]. Encounters [Hor09]. Encryption [Lan00a, Lan00b]. encyclopedia [Slo03]. end [Eps06]. energy [HS04]. engineering [Qua09]. Enigma [Cas06b]. enormous [WAD09]. equation [Len02]. equivalence [Dev04a]. Eratosthenes [Cox08]. Ernst [Moo09]. escape [Daw02]. Escher [dSLDS03]. Essay [Ano01g]. Essential [Hod06]. Euclid [Cas07, Gra08, Har00]. evaluation [Mol02]. Evansville [Was08]. Everyman [Bor03]. Everything [Har04]. Evolution [Grä09b]. Ewing [Jac09b]. examples [BB05]. exceptional [Agr08]. executive [Jac09a]. exhibition [Sig06]. expander [Sar04]. Expansions [Ise02]. Experiment [Sha05]. Experimental [BB05, McG01]. Experimentation [Sha05]. exploring [PC05].

F [Har04]. faces [Dur09]. facial [DWZ06]. factorizations [Jor03]. Faddeev [Ano08a]. Faisal [Jac06c]. family [HB09, Lyu00, Was08]. fascinating [AB08]. father [Bat08]. Fatou [Dev04b]. Feit [SST^{+}05]. Felix [CT07]. Fermat [Per00]. Feynman [Mar06]. FFT [MR01]. Fields [Ano06a, Mil08b, Rog00, FRS03, MP07, Rie02, VBKR07]. fiftieth [Ale07]. finance [App04]. Find [Lan06]. finds [Jac07a]. finite [Asc04]. first [Kra07, Was08, Lan00a]. flip [Cor04]. flow [And04]. Flowers [Ada05]. flows [Bra02]. fluctuations [Wee04]. fluid [Khe05]. forcing [Jec08].

Formal [Gon08, Hal08, Har08, Wie08]. Formatting [Grä09c]. forming [Jac06b]. formulas [AB08]. foundations [BC06]. four [DGPV07, Gon08]. four-color [Gon08]. Fourier [FS05, Ios01]. Francesco [BG09]. Franklin [BCG^{+}01]. Free [Fri00, Grä09a, CK09, DS04, Jac00f]. frequency [CHW09]. friend [Wel07]. Frogs [Dys09]. function [Boo08]. functional [Kad08]. functions [Bon02, Rue02]. Fundamental [KN08]. funny [Har04]. Furstenberg [Ano07i]. future [Bro02, Jac00b, Jac06d].

G [Ada05, KS04, RS09]. Galois

[Kis07b, Mil08b]. game [Lan04]. Gardner [Ano05b]. Gauss [Jac06a]. gear [Jac06d]. General [Per00, Bra02]. generalized [KTT09, FS05]. generate [Mez07]. Geneva [Sha09]. genomics [Kar02]. genus [HK09, Och09]. geometer [Ber00a, Ber00b]. geometric [Bra02]. Geometrization [And04, Jac06b]. Geometry [TU00, Aud08, Har00, Mar06, Sem03b, Tre00, Ser08]. George [Bar01, Ano09k, Bar01, Cas08, CJW07, DR07]. getting [Wie08]. Gian [BBS^{+}00]. Gian-Carlo [BBS^{+}00]. gift [Ono06]. Girgensohn [Sha05]. Gleason [Ano09h]. Glimm [Jac09c]. Gödel [Daw02, FK06, Fra06, Daw06, Dav01]. got [Kop00]. graduate [Fri05]. grand [Jaf06]. graph [Moh06]. Grauert [Huc09]. Gravity [Pet10]. great [Abh02, Gra08, WAD09, Zie04]. greatest [Szp03]. Gröbner [Stu05]. Gromov [Ano09j, Jac04c]. Grothendieck [Jac04b, Jac04a, Jac08a, Sch08]. Ground [PC05]. group [Agr08, Art00, Kis07a, Maj06, MR01, Sha01, Wri04, Aus09]. groups [App04, Asc04, Grü06, Gun06, Mez07, SC01, Sem03b]. grow [Ise02]. Guido [BG09].

H [Ano02a, Ano07b, CCJ^{+}01]. H.

[EGMW03, Spr00]. Hale [Mor05a]. Halmos [Ewi07]. Hans [Huc09]. Happy [Jac04d]. hard [Ree04]. Hardy [RMM08]. Harmonic [Art00]. Harmonious [KN08]. Harvard [Bat09]. Hash [Lan06]. Heath [Gra08]. heavens [Oss05]. Heiberg [Gra08]. Heilbronn [Sto06]. Heinz [Jac00e]. Heisenberg [Sem03b]. Heisuke [Jac05c]. helped [Szp03]. Henri [Aud09, DM08, Jac04d, Maw05]. Henriksen [DLM^{+}07]. heritage [dSLDS03]. Hermann [Eck06]. Hermitian [KT01]. Higgs [BGPG07]. high [Jac06d]. Hilbert [GG00, Its03]. Hironaka [Jac05c]. history [Rie02, Row01, Szp03]. hit [AB08]. Hofer [Ano00d]. holes [Bra02]. Homage [Aud09]. home [Jac07a]. Homological [dSG07]. Honeycombs [Hal00, KT01]. Honoring [Ono06]. honors [Was08]. horseshoe [Shu05]. Houghton [Sha09]. Hui [Cul02]. Hyman [Jac01c, Jac03c]. hyperbolic [Cou06]. Hyperion [Erm04]. hyperquadric [Tre00]. hypothesis [Con03, Woo01a, Woo01b, Woo02, Boo06].

IAS [FG05, Jac02a]. ice [Ada05, Gol09a]. ICM2006 [Jac06d]. icosahedron [CK05]. idea [DS01]. if [Jac04a, Jac04b, Cas07]. IHES [Jac08a]. II [Ber00b, Elk00b, Grä09b, Jac04b, Woo01b, Woo02]. III [Grä09c]. image [CSV03]. imaging [FS05]. impact [Fef06a, Fra06]. Implementations [Grä09b]. implications [BB05]. importance [Min05]. incompleteness [Dav06a, Fef06a, Fra06]. inequality [Bra02]. infinite [Poé07]. infinity [Har04, WBD^{+}04]. influence [Bat07]. Information [Ano01g]. insight [Bat07]. Institute [Ale07, Jac00c, Jac04h, Jac05a, Jac05a]. institutes [Jac09d]. integer [Slo03]. Integrable [Dei00, Its03]. integrals [Mol02]. intelligence [Gro00]. interactions [MS05]. interdependence [MS05]. Internet [Bis04, WAD09]. Interview [And07, Ano05b, Ano07j, Ano07k, Cas08, Jac01a, Jac01b, Jac02b, Jac04e, Jac05c, Jac08b, Per05, RS04, RS05, RS06, RS07, RS08, RS09, Szp07, Ano05d, Ano09k, Gel09, Jac01c, Jac03c, Jac03b, Jac09c, Jac05d]. Interviews [MP07]. Intimations [WBD^{+}04]. introduction [Sem03a, Sem03b]. Invariant [IK00, KTT09, SC01, Sch09]. invariants [Elk00a, Elk00b]. invented [Cho06]. Invention [Coo06]. invisible [Goe06]. Irving [BL07, Kad08]. Isadore [RS05]. ISBN [Sha05]. Itô [Pro07].

J [Max08]. J. [Dav01]. Jacobson [BKM^{+}00]. Jacques [LMT01, RS09]. Jacques-Louis [LMT01]. James [Jac05d, Jac09c, Sil08, Tol08]. Jan [DTRB01]. Japan [Jac04h]. Japanese [Cas07]. Jean [BHL00, RS04]. Jean-Pierre [RS04]. Jerry [Ano07j]. Joan [Ano07k]. John [Ano09m, Bri02, Jac09b, RS09]. Jon [GHW07]. Jonathan [Sha05]. Joseph [Jac04e]. journal [Jac00f]. JPEG [Aus08]. Jr [Dav01]. Julia [Dev04b]. Julian [BCF^{+}00]. Jürgen [MMNR00].

K. [Ada05]. Kantor [Abh02]. Kaplansky [BL07, Kad08]. Karp [Ano08c]. Kato [CJK^{+}00]. Keith [Bar01]. Keller [Jac04e]. Kepler [Szp03, Mor05a]. Kervaire [EdlHHW08]. kind [Gra03, Jac05a]. King [Jac06c]. Kissing [PZ04, Cas04]. Klee [GS08]. Klein [CT07]. Kleinberg [GHW07]. Kloosterman [Spr00]. Knew [Coo06]. knot [Sab09]. Krieger [Lan05]. Kumbakonam [Ono06]. Kurt [Dav01, Daw02, Daw06]. Kyosi [Pro07]. Kyoto [Ano06c, Ano08c].

L [Dav01, GS08]. labyrinth [Joh09]. Ladyzhenskaya [FLM^{+}04]. Lang [JK06, JK07]. Langlands [Ano07l]. last [Sil08]. lattice [Grä07]. Lattices [Elk00a, Elk00b]. Laurent [TPY03]. law [Rog00]. Lax [Ano05c, RS06, Sau00].

Learning [Cul02, PS03]. lecture [JW06, Dys09]. Lefschetz [Gom05]. Legendre [Dur09]. Legendrian [Sab09]. Lehmer [Bri07, Hir09]. length [Cal08]. Lennart [RS07]. Leonard [Hay01]. Leopold [Rei02]. Leray [BHL00]. Letter [DLM^{+}07, Kri05, Lan05]. letters [BG09]. LeVeque [Max08]. Lévy [App04]. Libbrecht [Ada05]. librarian [Rut02]. library [Jac07a, Rut02]. Lichnerowicz [KS09]. life [Dav01, Jac04a, Jac04b, Maw05, Moo09, Tu06]. Lifting [KS05]. Light [Pet10]. limits [Sla02]. line [Slo03]. Linear [Gal07, Elk00a, Elk00b]. Linking [Gro00]. Lions [LMT01]. Liu [Cul02]. local [Rog00]. logic [Dav01]. Logical [Dav01]. look [GG00]. Looking [Row01]. looks [Gra03, Jac06d]. lost [AB08, Gra08]. Louis [Jac02b, LMT01]. love [Coe09]. Luce [Ano05a]. lunch [DS04].

M [Ano09h, EGMW03]. M. [dSLDS03]. MA [Dav01]. machine [Cas06b, Cas06a]. Mackey [DR07]. Madhu [Wig03]. Man [Coo06]. Manifolds [Mor05b, And04, HS04, Mil03]. Manin [Gel09]. mapping [Hei06]. Margaret [Ano02a]. Markovic [DLM^{+}07]. Martin [Ano05b, Jac08b]. Math [Tan06, Szp03]. Mathematical [Cas06b, JS07, Kar02, Neu01, Qua09, Aus09, BCLP03, CMP04, Fri08, FRS03, GHW07, Hod01, JK07, Ree04, Ser08, Ste03, VBKR07]. Mathematician [Ano01g, Gra03]. Mathematicians [Rut02, Bar01, Jac02c, Kri04]. Mathematics [Dav05, DWZ06, Jac00b, Jac04f, MS05, Oss05, Saa08, Sau03, WAD09, BB05, Bat08, Bat09, Bro02, Cox05, Cul02, Fef06a, Fow05, Fri05, Gel09, Goe06, Gol09a, Gro00, Gro05, Jac04i, Jac06d, Jaf06, Kir02, Kob07, Kri05, Lub04, McC07, Min05, Pet09, PS03, Rut02, Sch07, Sus09, Was08, dSLDS03, Ale07, FG05, Jac02a, Jac05a, PA05, Sha05, Sas09, Sha05, BDK^{+}09, Coe09, Row01]. Mathematiker [Huc09]. MathML [Min05]. matrices [KT01, Mez07]. Matrix [Jor03, Dia05]. Matsusaka [Kol06]. Max [Daw02, Lax03]. Maxwell [Sil08]. McClure [Jac09a]. McNerney [Ano07j]. Me [Lan06]. measure [Ein09]. measuring [Abh02]. Medal [Ano05a, Jac06e, Jac07b, Rie02]. medalists [FRS03, MP07, VBKR07]. Medals [Ano06a]. Media [Gra03]. meets [Blu04, Dev04b]. memoriam [Jac05b]. Memories [Ber07, Poé08]. Mentors [Ano01g]. meromorphic [Bon02]. metaphor [McC07]. method [Gal07]. methods [BB05]. metric [Sem03a]. metrics [IK00]. Michael [JW06, RS05]. Michel [EdlHHW08]. mid [Kad08]. mid-twentieth [Kad08]. Mifflin [Sha09]. might [Sun08]. Mikio [And07, Sch07]. Millennium [Gun09, Jaf06]. mind [Gro00]. minds [Szp03]. minimal [Kol07]. minimum [HS04]. minor [Moh06]. minority [Goe06]. Mischa [Hor09]. miss [Sun08]. missing [Wee04]. mistaken [Dur09]. MN [Ada05]. Möbius [AR08]. Model [Pil00, Kol07, Lyu00]. modeling [EE03]. models [CSV03, Qua09]. modern [Art09]. module [Lic09]. moduli [Dev04a, Vak03]. molecular [Kar02]. Montgomery [Fin05]. Moore [Ano04b, Ano07b]. Morgan [Ano01a, Ano07c, Ano09d]. morphism [Kot04]. Moser [MMNR00]. Moses [Jac00e]. most [AB08]. mother [Art09]. motion [Sla02]. motive [Maz04]. MR0919604 [Row01]. MR1429389 [Dav01]. MR2016410 [Har04]. MR2125268 [Lan05]. MR2246251 [Gun09]. MR2259005 [Sus09]. MR2273791 [Sas09]. MR2279945 [Joh09]. MR2301067 [Moo09]. MR2352432 [Bla09]. MR2423374 [Aus09]. MR2457519 [Art09]. MR2467561 [BDK^{+}09]. MR2490399 [Coe09]. MSRI [Jac03a, Tan06]. Much [Coo06]. Multiscale [EE03]. Mumford [Ano06d]. Mushroom [PL06]. musical [CHW09]. mystery [Wee04].

named [Jac09a]. Narasimhan [Jac06c]. NAS [Jac04i]. Nathan [BKM^{+}00]. National [Jac06e, Ano05a, Jac07b]. nature [JW06, Sun08]. Navier [CF03]. néant [Jac04a, Jac04b]. needles [Tao01]. Nemmers [Ano08b, Jac00a, Jac04c]. nervous [Kop00]. networks [POM09, dSG07]. Nevanlinna [Ano06b]. Newest [Jac00c]. Newton [Blu04]. Niels [PC05]. Nirenberg [Jac02b]. no [Jac06b]. Nobel [Ney06]. Noether [Art09]. nonabelian [Rog00]. noncommutative [Cun01]. Nonholonomic [BMZ05]. nonlinearization [FS05]. Norton [Har04]. note [Ser08]. notebook [AB08]. Notes [Sto06, Sau03]. Notices [Lan05]. NSF [Jac00c]. number [Ano02c, Hir09, Poo08, Zie04]. numbers [Gol09b, Maz00, PZ04]. numerati [Sha09].

Oberwolfach [Jac00d, Jac04g]. Old [Agr08]. oldest [Szp03]. Oleinik [JLM03]. Olga [FLM^{+}04, JLM03]. olympiads [Tan06]. on-line [Slo03]. one [Szp03]. Open [JS07]. opens [Jac00c]. operad [Sta04]. operators [KTT09, Rue02]. Optics [Per00]. orbits [Abh02]. order [Erm04]. Osgood [Bat09]. Ostrowski [Ano00d]. Oswald [Bat07]. other [Aus09]. our [Gel09]. outer [PC05, Vog08]. own [Ewi07].

P [BCG^{+}01, Bor03]. packing [Ste03]. packings [PZ04]. pages [Sha05]. palimpsest [Gra08]. Paper [Ano01g, Cro09, Kra07]. parade [AB08]. parity [FI09]. partitions [AO01]. past [Jac00b]. Path [KN08]. Patterns [Gro05]. Paul [Ewi07]. PDE [CSV03, Tao01]. Pell [Len02]. pencil [Cro09, Gom05]. pencil-and-paper [Cro09]. Penrose [Bra02]. people [Jac04f]. percolation [Kes06]. performance [Ber02]. period [CG08]. Perseus [Dav01]. personal [Mol02]. personalities [BG09]. perspective [Rut02]. Peter [RS06]. Peters [Dav01, Sha05]. Peterson [BCG^{+}01]. Ph.D [Was08]. phenomenon [FI09, Sto06]. Phillips [Sar00]. physical [Mar06]. physics [MS05, PC05]. Pictures [Cas00, Cou06, Sig06]. Pierre [RS04]. place [Sau03]. Playright [Ano01g]. poem [Sil08]. poems [Coe09]. Poincaré [Jac06b, Maw05, Mil03, Wee04]. points [HS04]. polynomials [HM07]. popular [Fra06]. portrait [Dur09]. possible [Sus09]. potential [WAD09]. powers [Maz00]. practice [Har08]. Prague [Ber07]. Pratt [DLM^{+}07]. present [Grü06]. President [Art05]. Presidential [Ano05d, Ano09k, Jac01c, Jac03c, Jac03b, Jac05d, Jac09c]. Press [Ada05, Erm04]. prime [Zie04]. PRIMES [Bor03]. Princeton [BDK^{+}09]. Principle [Per00]. Prize [Ano09b, Gun09, Ano00a, Ano00d, Ano00b, Ano01b, Ano01a, Ano01c, Ano01e, Ano01f, Ano02b, Ano02d, Ano02c, Ano04f, Ano04a, Ano04g, Ano04b, Ano04d, Ano04e, Ano05c, Ano06c, Ano06d, Ano06b, Ano07a, Ano07i, Ano07b, Ano07l, Ano07c, Ano07d, Ano07e, Ano07g, Ano07m, Ano07h, Ano08a, Ano08b, Ano08c, Ano09a, Ano09c, Ano09i, Ano09j, Ano09d, Ano09e, Ano09g, Jac00a, Jac03e, Jac03d, Jac04c, Jac05e, Jac06c, Jac06a, Ney06, RS07]. Prizes [Ano00c, Ano01d, Ano02e, Ano04c, Ano07f, Ano09f]. Probability [SC01, App04]. problem [Its03, Mon01]. problems [Fri00, GG00, Grä07, Szp03, Gun09]. processes [App04]. processing [CSV03]. product [KTT09]. profession [Gel09]. Program [Ano01g, Kur00]. programming [Gal07]. Project [PA05]. projective [Gol07]. Proof [Ano01g, Hal08, Gon08, Har08, Jac06b, Wie08, Mor05a]. proofs [Cas00, PZ04]. property [Lub05, NY08]. Protocols [CT07]. pseudoholomorphic [Don05]. Publ [Dav01]. Public [Ano02a]. publication [Gra08, Lan05]. publish [Pet09]. publisher [Hod01]. Pulitzer [Ano01g]. pur [Huc09]. puzzles [Cro09].

quadratic [Lyu00]. qualitatively [Lyu00]. Quantum [Cun01, Lan04, App04, Maj06, Rud08]. quasi [Kot04]. quasi-morphism [Kot04]. quasiconformal [Hei06]. quasicrystal [Sen06]. quest [Daw06, Lip09]. Questions [Maz00, Knu02]. Quiver [DW05].

Rabinowitz [DF07]. races [Zie04]. Ralph [Sar00]. Ramanujan [AB08]. random [Dia05, Mez07, SC01]. Raoul [Jac01b, Tu06]. ratio [Lab08]. rationalists [Myc06]. Ray [Per00]. real [Dev04a, Gin09]. reals [Blu04, BC06]. receive [Ano03, Ano04f, Ano05a, Ano06d, Ano07i, Ano08a, Ano09i, Jac03d, Jac06c]. receives [Ano04g, Ano05c, Ano06c, Ano07m, Ano08b, Ano08c, Ano09j, Jac00a, Jac00e, Jac03e, Jac04c, Jac04i, Jac05e, Jac06e, Jac07b]. recipient [RS07]. reciprocity [Rog00]. record [Zie04]. Reducibility [Dav06b]. References [Grä09c]. Reflections [Bro02, Ano05d, Daw06]. relationship [Kob07]. Relativity [Per00, Bra02]. remarkable [Fri08]. remembered [RMM08]. Remembering [Ano09l, Ano09m, DF07]. Reply [DLM^{+}07]. representation [Kis07b]. representations [Art00, DW05]. research [Gro05, Leo08]. resonance [Jac07a]. retires [Jac09b]. Reunion [Ano01g]. revealed [AR08]. Review [Coo06, Erm04, Hod06, Mor05a, Sha05, Ada05, Art09, Aus09, BDK^{+}09, Bla09, CMP04, Coe09, Dav01, Gol09b, Gun09, Har04, Joh09, Moo09, Sas09, Sha09, Sus09, Gra03]. Revolution [Grä09b, Jac00f]. rhythm [CHW09, Kop00]. Ricci [And04]. Riemann [Boo06, Con03, Its03]. rigidity [Ein09]. RIMS [Jac04h]. ring [Vak03]. Rivest [Ano03]. Robbins [Ano07d]. Robert [Jac00e, KS04]. Roland [Sha05]. role [Cox05, Kad08]. Romania [Sau03]. Ross [Ano01g, Jac01a, JS03]. Rota [BBS^{+}00, Guo09]. rotating [Tao01]. route [Daw02].

S [EGMW03]. S. [Gra03, Jac05b]. S.-S. [Jac05b]. Sato [And07, Jac03d, Sch07]. Satter [Ano01c, Ano07e, Ano09e]. Scaling [Sla02]. scholars [Gra08]. School [FG05, Jac02a]. Schwartz [TPY03]. Schwarzian [OT09]. science [Erm04, Fri00, Gra03, Maw05, Qua09, Ano05a, Jac06e, Jac07b, Gra03]. scientific [BC06]. sea [Gol09a]. search [RSW06]. secret [Ada05]. Security [Lan00a]. Selberg [Ano09l]. sense [Sas09]. sensor [dSG07]. sequences [Cho06, Slo03]. Serge [JK06, JK07]. series [CMP04]. Serre [Jac03e, RS04]. service [Maw05, Ano02a]. set [FK06, Myc06]. sets [Zit09]. Severi [BG09]. Shamir [Ano03]. Shannon [GBC^{+}02]. shaped [Grä07]. share [Ano00d]. Shaw [Ano06d, Ano08a, Ano09i, Jac05e, Ano04g, Ano07l]. Shiffman [Lax03]. Shourik [Poé08]. SIAM [Ano00a, Ano09a]. Siberian [Daw02]. sideways [GG00]. Sierpinski [Dev04b]. sieve [Cox08]. simple [Asc04, RSW06]. simplex [Gal07]. Singer [Ano04f, RS05]. Singular [Bon02]. Sir [JW06]. sixtieth [Jac04g]. skein [Lic09]. slow [Jac00f]. Smale [Ano07i, Szp07]. small [RMM08, Sau03]. smart [Gol09b]. snowflake [Ada05]. Sobolev [Ale07]. Software [JS07]. solitons [TU00]. solution [Mon01]. solvable [Lyu00]. solve [Szp03]. Solving [Len02, Cro09]. Some [Kri04, PZ04, Szp03]. sometimes [Har04]. Sought [Ano01g]. Source [JS07, WAD09]. space [JW06, PC05, Roe06, Vak03, Vog08, Wee04]. spaces [Cun01, Dev04a, Lip09, Sem03a]. Spanish [Jac06d]. spatial [Neu01]. spectral [Cho06]. Spencer [KGG^{+}04]. sphere [Gin09, PZ04]. spiral [CM03]. spontaneous [Erm04]. squares [HM07]. Srinivasa [RS08]. stability [Tao01]. stable [Cal08]. staircases [CM03]. Stallings [Ano09m]. Standard [Lan00a, Lan00b]. Standing [Lan00b]. Stanley [PFTY09]. started [Wie08]. Starting [Kis07a]. static [DS01]. status [Asc04]. Steele [Ano00c, Ano01d, Ano02e, Ano04c, Ano07f, Ano09f]. Stephen [Szp07]. Steve [Erm04]. Stillwater [Ada05]. story [Mol02, Sti01]. strange [Rue06, Coe09]. Strasbourg [Aud08]. Strogatz [Erm04]. strong [CK09]. structure [Gol07]. Struik [DTRB01, Row01]. Stubhaug [Per05]. students [Fri05]. subdegrees [Abh02]. Sudan [Wig03]. Sudoku [HM07, Cro09]. Sullivan [Jac06e]. summoned [Jac04a, Jac04b]. sums [KT01]. super [Sla02, Gol09b]. super-Brownian [Sla02]. supposedly [Har04]. surfaces [Bon02]. surgery [DWZ06]. Survey [PA05]. swindle [Poé07]. symmetric [RSW06]. symmetry [Ada05, Grü06]. Sync [Erm04]. system [Kop00, Myc06]. systems [Cou06, Dei00, Its03, Kop00]. systole [Ber08]. syzygy [Wie06].

T [HO04]. table [Vog07]. tableau [Yon07]. tale [Ste03]. Tarski [Wel07]. Tate [Jac03d]. Taubes [Ano09i]. tautological [Vak03]. Taylor [Ano07l]. teach [Fri05]. Teaching [Fri05, Har00]. technology [Fri00]. ten [AB08]. tension [Sas09]. tensor [KTT09]. Teruhisa [Kol06]. test [Lan00b]. texts [Fri08]. their [Cun01, FS05]. Theorem [KN08, CK09, Dav06a, DGPV07, Fra06, Gon08]. theorems [Abh02, Fef06a]. theory [Ano02c, Aus09, Cas06b, Dei00, FK06, Grä07, Har08, KTT09, Lan04, Lip09, MR01, Mil08b, Myc06, Pil00, Poo08, Sha01]. Thesis [Fef06b]. thinking [Gol09b]. thinking-by-numbers [Gol09b]. Thomas [CCJ^{+}01]. Thompson [RS09]. Three [Abh02, Cas04, GG00, Mon01, MP07]. three-body [Mon01]. tiles [Aus05]. Time [CHW09, DS01, Jac00b, Lan00b]. Time-frequency [CHW09]. Tits [RS09]. today [Jac00d]. Too [Coo06]. top [AB08]. Topological [Khe05]. topology [Cun01, Dev04b, HK09]. topos [Ill04]. toric [Mil08a]. Tosio [CJK^{+}00]. trans [Daw02]. trans-Siberian [Daw02]. transfer [Rue02]. transform [Ios01]. transformations [AR08]. transformed [FK06]. transforms [FS05]. translation [Gra08, Lan05]. transport [Gol09a]. treasure [Tre00]. trees [Ise02]. Tribute [KS09, Fin05]. trio [Jac09d]. tropical [Mik07]. trove [Tre00]. Tukey [MR01, Bri02]. Turing [Ano03, Coo06, Blu04, Boo06, Cas06a, Dav06b, Fef06b, Hod06]. Tutte [HO04]. twentieth [Jac03a, Kad08]. twenty [GG00, Lan00a]. Twenty-first [Lan00a]. twenty-three [GG00]. Two [Grä07, BG09, Gra08].

U.S. [Tan06]. Uncovering [Boo08].

Undecidability [Poo08]. Undergraduate [Ano01g, Leo08]. uneasy [Kob07]. unexpected [PZ04]. universal [Lip09]. universe [DS01]. University [Jac07a].

value [BCLP03]. Varadhan

[Ano07m, RS08]. Variational [CSV03].

variety [Mil08a]. Veblen

[Ano01e, Ano04d, Ano07g, Bat07]. Venn [RSW06]. verification [Kur00]. vertex [DGPV07]. via [And04, HS04]. Victor [GS08]. Vietoris [Rei02]. views [Ano09k, Jac01c, Jac03c, Jac03b, Jac05d, Jac09c]. vision [Bat07]. visionary [Sch07]. Visualizing [Cox08]. Voiculescu [Jac04i]. void [Jac04a, Jac04b]. Volunteers [Ano01g]. voting [Saa08]. Voyageur [Ada05].

W [Dav01, Har04]. W. [Dav01]. walks [SC01]. Wallace [Har04]. Walter [SST^{+}05]. want [Hod01]. wars [Bla09]. wavelets [Jor03]. waves [Tao01]. way [Cul02, Gol09b]. ways [Abh02, BJK06]. Weil [Kri05, Lan05]. Weisfeiler [Lub04]. Wellesley [Dav01]. Weyl [Eck06]. where [Blu04]. Whiteman [Ano01f, Ano09g]. Whither [Dav05]. Who [Coo06, Sch08]. whose [DS01]. Wiener [Ano00a, Ano04e, Ano07h]. Wilder [Bri02]. Wiles [Jac05e]. will [CK09]. William [HO04, Max08]. Wins [Ano01g]. winter [Ada05]. Witten [Jac00a]. Wolf [Ano07i, Jac03d]. Wolff [CCJ^{+}01]. Wolfram [Gra03]. Woodin [Ste07]. words [Ewi07]. work [Dav01, FRS03, GHW07, Pro07, Ser08, Spr00, VBKR07, Wig03, Moo09]. works [Tu06]. world [Jac02c, Jac04h, Szp03]. worlds [Sus09]. worm [Boa03]. Wright [Ano02a]. write [Kra07]. Wu [Ano06d].

year [Leo08]. yesterday [Jac00d]. York [Har04, Erm04, Row01]. Young [Yon07]. Yuri [Gel09].

Zermelo [Moo09]. zeta [Rue02]. Zurich [Eck06].

References

Andrews:2008:YHP

[AB08] George E. Andrews and Bruce C.

Berndt. Your hit parade: the top ten most fascinating formulas in Ramanujan’s lost notebook. Notices of the American Mathematical Society, 55(1):18– 30, January 2008. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Arthur:2004:AB

[ABC^{+}04] James Arthur, Enrico Bombieri, Komaravolu Chandrasekharan, Friedrich Hirzebruch, Gopal Prasad, Jean-Pierre Serre, Tonny A. Springer, and Jacques Tits. Armand Borel (1923– 2003). Notices of the American Mathematical Society, 51 (5):498–524, May 2004. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Abhyankar:2002:TWM

[Abh02] Shreeram S. Abhyankar. Three ways of measuring distance, three orbits, three subdegrees, or the great theorems of Cameron and Kantor. Notices of the American Mathematical Society, 49(7):759–769, August 2002. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Adam:2005:FIB

[Ada05]  John A. Adam.  Flowers

of ice — beauty, symmetry, and complexity: a review of The snowflake: winter’s secret beauty [Voyageur Press, Stillwater, MN, 2003] by K. G. Libbrecht. Notices of the American Mathematical Society, 52 (4):402–416, April 2005. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Agricola:2008:ONE

[Agr08] Ilka Agricola. Old and new on the exceptional group G_{2}. Notices of the American Mathematical Society, 55(8):922–929, September 2008. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Alexandrov:2007:SIM

[Ale07] Victor Alexandrov. Sobolev Institute of Mathematics celebrates its fiftieth anniversary. Notices of the American Mathematical Society, 54(11):1512– 1514, December 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anderson:2004:GMR

[And04] Michael T. Anderson. Geometrization of 3-manifolds via the Ricci flow. Notices of the American Mathematical Society, 51(2):184–193, February 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Andronikof:2007:IMS

[And07] Emmanuel Andronikof. Interview with Mikio Sato. Notices of the American Mathematical Society, 54(2):208–222, February 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2000:ASW

[Ano00a] Anonymous. 2000 AMS–SIAM Wiener Prize. Notices of the American Mathematical Society, 47(4):483–484, April 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Awarded to Alexandre J. Chorin and Arthur T. Winfree.

Anonymous:2000:CP

[Ano00b] Anonymous. 2000 Cole Prize. Notices of the American Mathematical Society, 47(4):481–482, April 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Awarded to Andrei Suslin and Aise Johan de Jong.

Anonymous:2000:SP

[Ano00c] Anonymous. 2000 Steele Prizes. Notices of the American Mathematical Society, 47(4):477–480, April 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2000:BHS

[Ano00d]  Anonymous.  Beilinson and

Hofer share Ostrowski Prize. Notices of the American Mathematical Society, 47(8):885, September 2000. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2001:MP

[Ano01a]  Anonymous.  2000 Morgan

Prize. Notices of the American Mathematical Society, 48 (4):413–414, April 2001. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2001:CP

[Ano01b] Anonymous. 2001 Conant Prize.

Notices of the American Mathematical Society, 48(4):418–419, April 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2001:SPb

[Ano01c] Anonymous. 2001 Satter Prize.

Notices of the American Mathematical Society, 48(4):411–412, April 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2001:SPa

[Ano01d] Anonymous. 2001 Steele Prizes.

Notices of the American Mathematical Society, 48(4):404–407, April 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2001:VP

[Ano01e] Anonymous. 2001 Veblen Prize.

Notices of the American Mathematical Society, 48(4):408–410, April 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2001:WP

[Ano01f] Anonymous. 2001 Whiteman Prize. Notices of the American Mathematical Society, 48 (4):416–417, April 2001. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2001:DYI

[Ano01g]  Anonymous.  Departments:

For your information: Proof playright wins Pulitzer; call for mathematician mentors; 45th reunion conference for Ross program; undergraduate paper competitions in Cryptologia; volunteers sought for AWM essay contest; corrections. Notices of the American Mathematical Society, 48(6):607–608, June 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2002:ADP

[Ano02a] Anonymous. 2002 Award for Distinguished Public Service [to Margaret H. Wright]. Notices of the American Mathematical Society, 49(4):479–480, April 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2002:BP

[Ano02b] Anonymous. 2002 Bôcher Prize. Notices of the American Mathematical Society, 49(4):472–475, April 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2002:CPN

[Ano02c] Anonymous. 2002 Cole Prize in number theory. Notices of the American Mathematical Society, 49(4):476–478, April 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2002:CP

[Ano02d] Anonymous. 2002 Conant Prize.

Notices of the American Mathematical Society, 49(4):481–482, April 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2002:SP

[Ano02e] Anonymous. 2002 Steele Prizes.

Notices of the American Mathematical Society, 49(4):466–471, April 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2003:RSA

[Ano03] Anonymous. Rivest, Shamir, and Adleman receive 2002 Turing Award. Notices of the American Mathematical Society, 50 (7):800, August 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2004:CP

[Ano04a] Anonymous. 2004 Conant Prize.

Notices of the American Mathematical Society, 51(4):433–434, April 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2004:MP

[Ano04b] Anonymous. 2004 Moore Prize.

Notices of the American Mathematical Society, 51(4):431–432, April 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2004:SP

[Ano04c] Anonymous. 2004 Steele Prizes. Notices of the American Mathematical Society, 51(4):421–425, April 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2004:VP

[Ano04d] Anonymous. 2004 Veblen Prize. Notices of the American Mathematical Society, 51(4):426–427, April 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2004:WP

[Ano04e] Anonymous. 2004 Wiener Prize. Notices of the American Mathematical Society, 51(4):428–430, April 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2004:ASR

[Ano04f] Anonymous. Atiyah and Singer receive 2004 Abel Prize. Notices of the American Mathematical Society, 51(6):649–650, June/July 2004. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2004:CRS

[Ano04g] Anonymous. Chern receives Shaw Prize. Notices of the American Mathematical Society, 51(9):1067–1068, October 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2005:BLR

[Ano05a] Anonymous. de Boor and Luce receive National Medal of Science. Notices of the American Mathematical Society, 52 (6):634, June/July 2005. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2005:IMG

[Ano05b] Anonymous. Interview with Martin Gardner. Notices of the American Mathematical Society, 52(6):602–611, June/July 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2005:LRA

[Ano05c] Anonymous. Lax receives Abel Prize. Notices of the American Mathematical Society, 52 (6):632–633, June/July 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2005:PRI

[Ano05d] Anonymous. Presidential reflections: interview with David Eisenbud. Notices of the American Mathematical Society, 52 (2):216–218, February 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2006:FMA

[Ano06a] Anonymous. 2006 Fields Medals awarded. Notices of the American Mathematical Society, 53 (9):1037–1044, October 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2006:NPA

[Ano06b] Anonymous. 2006 Nevanlinna Prize awarded. Notices of the American Mathematical Society, 53(9):1045–1047, October 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2006:ARK

[Ano06c] Anonymous. Akaike receives 2006 Kyoto Prize. Notices of the American Mathematical Society, 53(9):1056–1057, October 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2006:MWR

[Ano06d] Anonymous. Mumford and Wu receive 2006 Shaw Prize. Notices of the American Mathematical Society, 53(9):1054–1055, October 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2007:CP

[Ano07a] Anonymous. 2007 Conant Prize. Notices of the American Mathematical Society, 54(4):519–520, April 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2007:HMP

[Ano07b] Anonymous. 2007 E. H. Moore Prize. Notices of the American Mathematical Society, 54 (4):532–533, April 2007. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2007:MP

[Ano07c]  Anonymous.  2007 Morgan

Prize. Notices of the American Mathematical Society, 54 (4):521–522, April 2007. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2007:RP

[Ano07d]  Anonymous.  2007 Robbins

Prize. Notices of the American Mathematical Society, 54 (4):534–535, April 2007. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2007:SPb

[Ano07e] Anonymous. 2007 Satter Prize.

Notices of the American Mathematical Society, 54(4):523–524, April 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2007:SPa

[Ano07f] Anonymous. 2007 Steele Prizes.

Notices of the American Mathematical Society, 54(4):514–518, April 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2007:VP

[Ano07g] Anonymous. 2007 Veblen Prize.

Notices of the American Mathematical Society, 54(4):527–530, April 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2007:WP

[Ano07h] Anonymous. 2007 Wiener Prize. Notices of the American Mathematical Society, 54(4):525–526, April 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2007:FSR

[Ano07i] Anonymous. Furstenberg and Smale receive 2006–2007 Wolf Prize. Notices of the American Mathematical Society, 54 (5):631–632, May 2007. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2007:ICJ

[Ano07j] Anonymous. Interview with Congressman Jerry McNerney. Notices of the American Mathematical Society, 54(9):1158– 1160, October 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2007:IJB

[Ano07k] Anonymous. Interview with Joan Birman. Notices of the American Mathematical Society, 54(1):20–29, January 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2007:LTA

[Ano07l] Anonymous. Langlands and Taylor awarded Shaw Prize. Notices of the American Mathematical Society, 54(8):1001, September 2007. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2007:VRA

[Ano07m] Anonymous. Varadhan receives 2007 Abel Prize. Notices of the American Mathematical Society, 54(6):738–739, June/July 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2008:AFR

[Ano08a] Anonymous. Arnold and Faddeev receive 2009 Shaw Prize. Notices of the American Mathematical Society, 55(8):966, September 2008. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2008:DRN

[Ano08b] Anonymous. Donaldson receives Nemmers Prize. Notices of the American Mathematical Society, 55(7):808, August 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2008:KRK

[Ano08c]  Anonymous.  Karp receives

2009 Kyoto Prize. Notices of the American Mathematical Society, 55(9):1102–1103, October 2008. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2009:ASB

[Ano09a] Anonymous. 2009 AMS–SIAM Birkhoff Prize. Notices of the American Mathematical Society, 56(4):492–493, April 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2009:CPA

[Ano09b] Anonymous. 2009 Cole Prize in Algebra. Notices of the American Mathematical Society, 56 (4):494–495, April 2009. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2009:CP

[Ano09c] Anonymous. 2009 Conant Prize. Notices of the American Mathematical Society, 56(4):500–501, April 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2009:MP

[Ano09d]  Anonymous.  2009 Morgan

Prize. Notices of the American Mathematical Society, 56 (4):502–503, April 2009. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2009:SPb

[Ano09e] Anonymous. 2009 Satter Prize. Notices of the American Mathematical Society, 56(4):496, April 2009. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2009:SPa

[Ano09f] Anonymous. 2009 Steele Prizes.

Notices of the American Mathematical Society, 56(4):488–491, April 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2009:WP

[Ano09g] Anonymous. 2009 Whiteman Prize. Notices of the American Mathematical Society, 56 (4):498–499, April 2009. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2009:AMG

[Ano09h] Anonymous. Andrew M. Gleason 1921–2008. Notices of the American Mathematical Society, 56(10):1236–1267, November 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Ethan D. Bolker, coordinating editor.

Anonymous:2009:DTR

[Ano09i] Anonymous. Donaldson and Taubes receive 2009 Shaw Prize. Notices of the American Mathematical Society, 56(8):952–953, September 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2009:GRA

[Ano09j] Anonymous. Gromov receives 2009 Abel Prize. Notices of the American Mathematical Society, 56(6):730–731, June/July 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:2009:PVI

[Ano09k] Anonymous. Presidential views: interview with George Andrews. Notices of the American Mathematical Society, 56(3):383–385, March 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Conducted by Allyn Jackson.

Anonymous:2009:RAS

[Ano09l] Anonymous. Remembering Atle Selberg, 1917–2007. Notices of the American Mathematical Society, 56(6):692–710, June/ July 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Dennis Hejhal, coordinating editor.

Anonymous:2009:RJS

[Ano09m] Anonymous. Remembering John Stallings. Notices of the American Mathematical Society, 56(11):1410–1417, December 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Danny Calegari and Benson Farb, coordinating editors.

Ahlgren:2001:ACA

[AO01] Scott Ahlgren and Ken Ono. Addition and counting: the arithmetic of partitions. Notices of the American Mathematical Society, 48(9):978–984, October 2001. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Applebaum:2004:LPP

[App04] David Applebaum. Lévy processes — from probability to finance and quantum groups. Notices of the American Mathematical Society, 51(11):1336– 1347, December 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Arnold:2008:MTR

[AR08] Douglas N. Arnold and Jonathan Rogness. Möbius transformations revealed. Notices of the American Mathematical Society, 55(10):1226–1231, November 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Arthur:2000:HAG

[Art00] James Arthur. Harmonic analysis and group representations. Notices of the American Mathematical Society, 47(1):26–34, January 2000. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Arthur:2005:APA

[Art05] James Arthur. AMS President’s address at Abel celebration. Notices of the American Mathematical Society, 52(7):745–746, August 2005. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Artmann:2009:ENM

[Art09]  Benno Artmann.  Emmy

Noether: the mother of modern algebra [book review of MR2457519]. Notices of the American Mathematical Society, 56(9):1105, October 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Aschbacher:2004:SCF

[Asc04] Michael Aschbacher. The status of the classification of the finite simple groups. Notices of the American Mathematical Society, 51(7):736–740, August 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Audin:2008:DGS

[Aud08] Michèle Audin. Differential geometry, Strasbourg, 1953. Notices of the American Mathematical Society, 55(3):366–370, March 2008. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Audin:2009:HHC

[Aud09] Michèle Audin. Homage to Henri Cartan (1904–2008). Notices of the American Mathematical Society, 56(5):614–615, May 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Austin:2005:T

[Aus05] David Austin. Up and down the tiles. Notices of the American Mathematical Society, 52 (6):600–601, June/July 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Austin:2008:WLJ

[Aus08] David Austin. What is … JPEG? Notices of the American Mathematical Society, 55(2): 226–229, February 2008. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Austin:2009:GTB

[Aus09] David Austin. Group theory in the bedroom, and other mathematical diversions [book review of MR2423374]. Notices of the American Mathematical Society, 56(2):237–239, February 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Barenblatt:2001:GKB

[Bar01]  G. I. Barenblatt.  George

Keith Batchelor (1920–2000) and David George Crighton (1942–2000): applied mathematicians. Notices of the American Mathematical Society, 48 (8):800–806, September 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Batterson:2007:VII

[Bat07] Steve Batterson. The vision, insight, and influence of Oswald Veblen. Notices of the American Mathematical Society, 54 (5):606–618, May 2007. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Batterson:2008:FFA

[Bat08] Steve Batterson. The father of the father of American mathematics. Notices of the American Mathematical Society, 55 (3):352–363, March 2008. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Batterson:2009:BOA

[Bat09] Steve Batterson. Bôcher, Osgood, and the ascendance of American mathematics at Harvard. Notices of the American Mathematical Society, 56 (8):916–928, September 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bailey:2005:EME

[BB05] David H. Bailey and Jonathan M. Borwein. Experimental mathematics: examples, methods and implications. Notices of the American Mathematical Society, 52(5):502–514, May 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

docserver.carma.newcastle.
edu.au/1241/;   http://www.
ams.org/notices/200505/fea-
borwein.pdf.

Beschler:2000:GCR

[BBS^{+}00] Edwin F. Beschler, David A.

Buchsbaum, Jacob T. Schwartz, Richard P. Stanley, Brian D. Taylor, and Michael Waterman. Gian-Carlo Rota (1932–1999). Notices of the American Mathematical Society, 47(2):203–216, February 2000. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Braverman:2006:CRF

[BC06] Mark Braverman and Stephen Cook. Computing over the reals: foundations for scientific computing. Notices of the American Mathematical Society, 53 (3):318–329, March 2006. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bluman:2000:JDC

[BCF^{+}00] George Bluman, L. Pamela Cook, Joe Flaherty, Jerry Kevorkian, Norman Malmuth, Robert O’Malley, Donald W. Schwendeman, and Marshall Tulin. Julian D. Cole (1925– 1999). Notices of the American Mathematical Society, 47 (4):466–473, April 2000. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Brown:2001:FPP

[BCG^{+}01] E. H. Brown, F. R. Cohen, F. W. Gehring, H. R. Miller, and B. A. Taylor. Franklin P. Peterson (1930–2000). Notices of the American Mathematical Society, 48(10):1161– 1168, November 2001. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Batterson:2003:VMA

[BCLP03] Steve Batterson, Charles Curtis, Albert Lewis, and Karen Parshall. The value of mathematical archives. Notices of the American Mathematical Society, 50(11):1409–1415, December 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Birch:2009:PCM

[BDK^{+}09] Bryan Birch, Simon Donaldson, Gil Kalai, Richard Kenyon, and Angus Macintyre. The Princeton companion to mathematics [book review of MR2467561]. Notices of the American Mathematical Society, 56(10):1276– 1281, November 2009. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Berger:2000:EG

[Ber00a] Marcel Berger. Encounter with a geometer. I. Notices of the American Mathematical Society, 47(2):183–194, February 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Berger:2000:EGI

[Ber00b] Marcel Berger. Encounter with a geometer. II. Notices of the American Mathematical Society, 47(3):326–340, March 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Berlekamp:2002:PBC

[Ber02] Elwyn Berlekamp. The performance of block codes. Notices of the American Mathematical Society, 49(1):17–22, January 2002. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bers:2007:MP

[Ber07] Lipman Bers. Memories of Prague. Notices of the American Mathematical Society, 54 (7):859–860, August 2007. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Berger:2008:WLS

[Ber08] Marcel Berger. What is … a systole? Notices of the American Mathematical Society, 55 (3):374–376, March 2008. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Babbitt:2009:GCF

[BG09] Donald Babbitt and Judith Goodstein. Guido Castelnuovo and Francesco Severi: two personalities, two letters. Notices of the American Mathematical Society, 56(7):800–808, August 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bradlow:2007:WLH

[BGPG07] Steven B. Bradlow, Oscar García-Prada, and Peter B. Gothen. What is … a Higgs bundle? Notices of the American Mathematical Society, 54 (8):980–981, September 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Borel:2000:JL

[BHL00] Armand Borel, Gennadi M. Henkin, and Peter D. Lax. Jean Leray (1906–1998). Notices of the American Mathematical Society, 47(3):350–359, March 2000. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Biss:2004:EI

[Bis04] Daniel K. Biss. The elephant in the Internet. Notices of the American Mathematical Society, 51(10):1217–1219, November 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Brams:2006:BWC

[BJK06] Steven J. Brams, Michael A. Jones, and Christian Klamler. Better ways to cut a cake. Notices of the American Mathematical Society, 53(11):1314– 1321, December 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Benkart:2000:NJ

[BKM^{+}00] Georgia Benkart, Irving Kaplansky, Kevin McCrimmon, David J. Saltman, and George B. Seligman. Nathan Jacobson (1910–1999). Notices of the American Mathematical Society, 47(9):1061–1071, October 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bass:2007:IK

[BL07] Hyman Bass and T. Y. Lam.

Irving Kaplansky 1917–2006. Notices of the American Mathematical Society, 54(11):1477– 1493, December 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Blank:2009:CWB

[Bla09] Brian E. Blank. The calculus wars [book review of MR2352432]. Notices of the American Mathematical Society, 56(5):602–610, May 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Blum:2004:CRW

[Blu04] Lenore Blum. Computing over the reals: where Turing meets Newton. Notices of the American Mathematical Society, 51 (9):1024–1034, October 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bloch:2005:ND

[BMZ05] Anthony M. Bloch, Jerrold E.

Marsden, and Dmitry V. Zenkov. Nonholonomic dynamics. Notices of the American Mathematical Society, 52 (3):324–333, March 2005. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Boas:2003:WLW

[Boa03] Harold P. Boas. What is … a worm? Notices of the American Mathematical Society, 50 (5):554–555, May 2003. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bonk:2002:SSM

[Bon02] Mario Bonk. Singular surfaces and meromorphic functions. Notices of the American Mathematical Society, 49 (6):647–657, June/July 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Booker:2006:TRH

[Boo06] Andrew R. Booker. Turing and the Riemann Hypothesis. Notices of the American Mathematical Society, 53(10):1208– 1211, November 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

                                        URL http://www.ams.org/
                                        notices/200610/fea-booker.
                                        pdf.

Booker:2008:UNF

[Boo08] Andrew R. Booker. Uncovering a new L-function. Notices of the American Mathematical Society, 55(9):1088–1094, October 2008. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Bornemann:2003:PPB

[Bor03] Folkmar Bornemann. PRIMES is in P: a breakthrough for “Everyman”. Notices of the American Mathematical Society, 50 (5):545–552, May 2003. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bray:2002:BHG

[Bra02] Hubert L. Bray. Black holes, geometric flows, and the Penrose inequality in general relativity. Notices of the American Mathematical Society, 49 (11):1372–1381, December 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Brillinger:2002:JWT

[Bri02]  David R. Brillinger.  John

Wilder Tukey (1915–2000). Notices of the American Mathematical Society, 49(2):193–201, February 2002. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Brillhart:2007:EL

[Bri07] John Brillhart. Emma Lehmer 1906–2007. Notices of the American Mathematical Society, 54 (11):1500–1501, December 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Browder:2002:RFM

[Bro02] Felix Browder. Reflections on the future of mathematics. Notices of the American Mathematical Society, 49(6):658–662, June/July 2002. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Calegari:2008:WLS

[Cal08] Danny Calegari. What is … stable commutator length? Notices of the American Mathematical Society, 55(9):1100– 1101, October 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Candes:2003:WLC

[Can03] Emmanuel J. Candès. What is … a curvelet? Notices of the American Mathematical Society, 50(11):1402–1403, December 2003. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Carlson:2005:CCC

[Car05] Jon F. Carlson. Cohomology, computations, and commutative algebra. Notices of the American Mathematical Society, 52 (4):426–434, April 2005. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Casselman:2000:PP

[Cas00] Bill Casselman. Pictures and proofs. Notices of the American Mathematical Society, 47(10): 1257–1266, November 2000.

CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Casselman:2004:DKT

[Cas04] Bill Casselman. The difficulties of kissing in three dimensions. Notices of the American Mathematical Society, 51(8):884–885, September 2004. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Casselman:2006:BTM

[Cas06a] Bill Casselman. Blueprint for a Turing Machine: About the cover… and a bit more. Notices of the American Mathematical Society, 53(10):1186– 1189, November 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/
notices/200610/comm-aboutcov.
pdf.

Casselman:2006:MTE

[Cas06b] Bill Casselman. Mathematical theory of the Enigma machine. Notices of the American Mathematical Society, 53(4):433, April 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://www.

mathaware.org/;   http://
www.nationalarchives.gov.
uk/;   http://www.turingarchive.
org/browse.php/C/30. The

front cover of this issue displays eight pages of Alan Turing’s description of the Enigma machine. The issue is a special tribute to Kurt Gödel for the centenary of his birth.

Casselman:2007:IEB

[Cas07] Bill Casselman. If Euclid had been Japanese. Notices of the American Mathematical Society, 54(5):626–628, May 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Casselman:2008:IGC

[Cas08] Bill Casselman. Interview with George Csicsery. Notices of the American Mathematical Society, 55(5):576–578, May 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Carleson:2001:THW

[CCJ^{+}01] Lennart Carleson, Sun-Yung Alice Chang, Peter W. Jones, Markus Keel, Peter D. Lax, Nikolai Makarov, Donald Sarason, Wilhelm Schlag, and Barry Simon. Thomas H. Wolff (1954– 2000). Notices of the American Mathematical Society, 48 (5):482–490, May 2001. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cannone:2003:NBC

[CF03] Marco Cannone and Susan Friedlander. Navier: blow-up and collapse. Notices of the American Mathematical Society, 50(1):7–13, January 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Carlson:2008:WLP

[CG08] James Carlson and Phillip Griffiths. What is … a period domain? Notices of the American Mathematical Society, 55 (11):1418–1419, December 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chow:2006:YCI

[Cho06]  Timothy Y. Chow.  You

could have invented spectral sequences. Notices of the American Mathematical Society, 53 (1):15–19, January 2006. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cheng:2009:TFA

[CHW09] Xiaowen Cheng, Jarod V. Hart, and James S. Walker. Timefrequency analysis of musical rhythm. Notices of the American Mathematical Society, 56 (3):356–372, March 2009. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cordes:2000:TK

[CJK^{+}00] Heinz Cordes, Arne Jensen, S. T. Kuroda, Gustavo Ponce, Barry Simon, and Michael Taylor. Tosio Kato (1917–1999). Notices of the American Mathematical Society, 47(6):650–657, June/July 2000. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cottle:2007:GBD

[CJW07] Richard Cottle, Ellis Johnson, and Roger Wets. George B. Dantzig (1914–2005). Notices of the American Mathematical Society, 54(3):344–362, March 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cornelissen:2005:AI

[CK05] Gunther Cornelissen and Fumiharu Kato. The p-adic icosahedron. Notices of the American Mathematical Society, 52 (7):720–727, August 2005. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Conway:2009:SFW

[CK09] John H. Conway and Simon Kochen. The strong free will theorem. Notices of the American Mathematical Society, 56(2):226–232, February 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Colding:2003:DDS

[CM03] Tobias H. Colding and William P. Minicozzi, II. Disks that are double spiral staircases. Notices of the American Mathematical Society, 50(3):327–339, March 2003. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Callaghan:2004:BCS

[CMP04] Thomas Callaghan, Peter J.

Mucha, and Mason A. Porter. The bowl championship series: a mathematical review. Notices of the American Mathematical Society, 51(8):887–893, September 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Coetzee:2009:SAP

[Coe09] J. M. Coetzee. Strange attractors: poems of love and mathematics [book review of MR2490399]. Notices of the American Mathematical Society, 56(8):944–946, September 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Conrey:2003:RH

[Con03] J. Brian Conrey. The Riemann hypothesis. Notices of the American Mathematical Society, 50(3):341–353, March 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cooper:2006:MWK

[Coo06] S. Barry Cooper. The Man Who Knew Too Much: Alan Turing and the Invention of the Computer — a book review. Notices of the American Mathematical Society, 53(10):1213– 1217, November 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/

                                        notices/200610/rev-cooper.
                                        pdf.

Corti:2004:WLF

[Cor04] Alessio Corti. What is … a flip? Notices of the American Mathematical Society, 51 (11):1350–1351, December 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Coudene:2006:PHD

[Cou06] Yves Coudene. Pictures of hyperbolic dynamical systems. Notices of the American Mathematical Society, 53(1):8–13, January 2006. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cox:2005:WRA

[Cox05] David A. Cox. What is the role of algebra in applied mathematics? Notices of the American Mathematical Society, 52(10):1193–1198, November 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cox:2008:VSE

[Cox08] David N. Cox. Visualizing the sieve of Eratosthenes. Notices of the American Mathematical Society, 55(5):579–582, May 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Crook:2009:PPA

[Cro09] J. F. Crook. A pencil-and-paper algorithm for solving Sudoku puzzles. Notices of the American Mathematical Society, 56 (4):460–468, April 2009. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chan:2003:VPM

[CSV03] Tony F. Chan, Jianhong Shen, and Luminita Vese. Variational PDE models in image processing. Notices of the American Mathematical Society, 50(1):14– 26, January 2003. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Chislenko:2007:FKP

[CT07] Eugene Chislenko and Yuri Tschinkel. The Felix Klein Protocols. Notices of the American Mathematical Society, 54 (8):960–970, September 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cullen:2002:LLH

[Cul02] Christopher Cullen. Learning from Liu Hui? A different way to do mathematics. Notices of the American Mathematical Society, 49(7):783–790, August 2002. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cuntz:2001:QST

[Cun01]  Joachim Cuntz.  Quantum

spaces and their noncommutative topology. Notices of the American Mathematical Society, 48(8):793–799, September 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Davis:2001:BRL

[Dav01] Martin Davis. Book review: Logical dilemmas: the life and work of Kurt Gödel [A. K. Peters, Wellesley, 1997; MR1429389 (98a:01017)] by J. W. Dawson, Jr., and Gödel: a life of logic [Perseus Publ., Cambridge, MA, 2000] by J. L. Casti and W. DePauli. Notices of the American Mathematical Society, 48(8):807–813, September 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Davies:2005:WM

[Dav05] Brian Davies. Whither mathematics? Notices of the American Mathematical Society, 52(11):1350–1356, December 2005. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

                                        URL http://www.ams.org/
                                        notices/200511/comm-davies.
                                        pdf.

Davis:2006:IT

[Dav06a] Martin Davis. The incompleteness theorem. Notices of the American Mathematical Society, 53(4):414–418, April 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Davis:2006:WLT

[Dav06b] Martin Davis. What is … Turing reducibility? Notices of the American Mathematical Society, 53(10):1218– 1219, November 2006. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL http://www.ams.
org/notices/200610/whatis-
davis.pdf.

Dawson:2002:MDK

[Daw02] John W. Dawson, Jr. Max Dehn, Kurt Gödel, and the trans-Siberian escape route. Notices of the American Mathematical Society, 49(9):1068– 1075, October 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dawson:2006:QKG

[Daw06] John W. Dawson, Jr. In quest of Kurt Gödel: reflections of a biographer. Notices of the American Mathematical Society, 53 (4):444–447, April 2006. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Deift:2000:ISC

[Dei00] Percy Deift. Integrable systems and combinatorial theory. Notices of the American Mathematical Society, 47(6):631–640, June/July 2000. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devadoss:2004:CER

[Dev04a] Satyan L. Devadoss. Combinatorial equivalence of real moduli spaces. Notices of the American Mathematical Society, 51 (6):620–628, June/July 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Devaney:2004:CSJ

[Dev04b] Robert L. Devaney. Cantor and Sierpinski, Julia and Fatou: complex topology meets complex dynamics. Notices of the American Mathematical Society, 51(1):9–15, January 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Davis:2007:RPR

[DF07] Philip J. Davis and Aviezri S. Fraenkel. Remembering Philip Rabinowitz. Notices of the American Mathematical Society, 54(11):1502–1506, December 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

DeTurck:2007:FVT

[DGPV07] Dennis DeTurck, Herman Gluck, Daniel Pomerleano, and David Shea Vick. The four vertex theorem and its converse. Notices of the American Mathematical Society, 54(2):192–207, February 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Diaconis:2005:WLR

[Dia05] Persi Diaconis. What is … a random matrix? Notices of the American Mathematical Society, 52(11):1348–1349, December 2005. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Daubechies:2007:LER

[DLM^{+}07] Ingrid Daubechies, Robert Lazarsfeld, John Morgan, Andrei Okounkov, and Terence Tao. Letter to the Editor: Reply to Davey, Henriksen, Markovic and Pratt. Notices of the American Mathematical Society, 54 (6):694, June/July 2007. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL http://www.ams.
org/notices/200706/200706-
full-issue.pdf.

Davis:2008:HCB

[DM08] Philip J. Davis and David Mumford. Henri’s crystal ball. Notices of the American Mathematical Society, 55(4):458–466, April 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Donaldson:2005:WLP

[Don05] Simon K. Donaldson. What is … a pseudoholomorphic curve? Notices of the American Mathematical Society, 52(9):1026– 1027, October 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Doran:2007:GM

[DR07] Robert S. Doran and Arlan Ramsay. George Mackey 1916– 2006. Notices of the American Mathematical Society, 54 (7):824–850, August 2007. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Daigneault:2001:ESU

[DS01] Aubert Daigneault and Arturo Sangalli. Einstein’s static universe: an idea whose time has come back? Notices of the American Mathematical Society, 48(1):9–16, January 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Delbaen:2004:WLF

[DS04] Freddy Delbaen and Walter Schachermayer. What is … a free lunch? Notices of the American Mathematical Society, 51(5):526–528, May 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

deSilva:2007:HSN

[dSG07] Vin de Silva and Robert Ghrist. Homological sensor networks. Notices of the American Mathematical Society, 54(1):10–17, January 2007. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

deSmit:2003:AMH

[dSLDS03] B. de Smit, H. W. Lenstra, Jr., Douglas Dunham, and Reza Sarhangi. Artful mathematics: the heritage of M. C. Escher. Notices of the American Mathematical Society, 50(4):446–457, April 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Davis:2001:DJS

[DTRB01] Chandler Davis, Jim Tattersall, Joan Richards, and Tom Banchoff. Dirk Jan Struik (1894– 2000). Notices of the American Mathematical Society, 48 (6):584–589, June/July 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Duren:2009:CFM

[Dur09] Peter Duren. Changing faces:

the mistaken portrait of Legendre. Notices of the American Mathematical Society, 56 (11):1440–1443, December 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Derksen:2005:QR

[DW05] Harm Derksen and Jerzy Weyman. Quiver representations. Notices of the American Mathematical Society, 52(2):200–206, February 2005. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Deuflhard:2006:MFS

[DWZ06] Peter Deuflhard, Martin Weiser, and Stefan Zachow. Mathematics in facial surgery. Notices of the American Mathematical Society, 53(9):1012–1016, October 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dyson:2009:BFE

[Dys09] Freeman J. Dyson. Birds and frogs: 2008 Einstein Lecture. Notices of the American Mathematical Society, 56(2):212– 223, February 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

                                        URL http://www.ams.org/
                                        notices/200902/rtx090200212p.
                                        pdf.

Eckmann:2006:HWZ

[Eck06] Beno Eckmann. Hermann Weyl in Zurich 1950–1955. Notices of the American Mathematical Society, 53(10):1222–1223, November 2006. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Eliahou:2008:MK

[EdlHHW08] Shalom Eliahou, Pierre de la Harpe, Jean-Claude Hausmann, and Claude Weber. Michel Kervaire 1927–2007. Notices of the American Mathematical Society, 55(8):960–961, September 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

E:2003:MMC

[EE03] Weinan E and Bjorn Engquist. Multiscale modeling and computation. Notices of the American Mathematical Society, 50 (9):1062–1070, October 2003.

CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ellers:2003:HMC

[EGMW03] Erich W. Ellers, Branko Grünbaum, Peter McMullen, and Asia Ivić Weiss. H. S. M. Coxeter (1907–2003). Notices of the American Mathematical Society, 50(10):1234–1240, November 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Einsiedler:2009:WLM

[Ein09] Manfred Einsiedler. What is … measure rigidity? Notices of the American Mathematical Society, 56(5):600–601, May 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Elkies:2000:LLCa

[Elk00a] Noam D. Elkies. Lattices, linear codes, and invariants. I. Notices of the American Mathematical Society, 47(10):1238– 1245, November 2000. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Elkies:2000:LLCb

[Elk00b] Noam D. Elkies. Lattices, linear codes, and invariants. II. Notices of the American Mathematical Society, 47(11):1382– 1391, December 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Epstein:2006:WLB

[Eps06] Charles L. Epstein. What is … a bad end? Notices of the American Mathematical Society, 53(9):1028–1029, October 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ermentrout:2004:RSE

[Erm04] G. Bard Ermentrout. Review of Sync: The emerging science of spontaneous order [Hyperion Press, New York, 2003] by Steve Strogatz. Notices of the American Mathematical Society, 51 (3):312–319, March 2004. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ewing:2007:PHH

[Ewi07] John Ewing. Paul Halmos: in his own words. Notices of the American Mathematical Society, 54(9):1136–1144, October 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Feferman:2006:IIT

[Fef06a] Solomon Feferman. The impact of the incompleteness theorems on mathematics. Notices of the American Mathematical Society, 53(4):434–439, April 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Feferman:2006:TT

[Fef06b]  Solomon Feferman.  Tur-

ing’s thesis. Notices of the American Mathematical Society, 53(10):1200–1206, November 2006. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/
notices/200610/fea-feferman.
pdf.

Friedlander:2005:ISM

[FG05] Susan Friedlander and Mark Goresky. The IAS School of Mathematics at 75. Notices of the American Mathematical Society, 52(8):859–862, September 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Friedlander:2009:WLP

[FI09] John Friedlander and Henryk Iwaniec. What is … the parity phenomenon? Notices of the American Mathematical Society, 56(7):817–818, August 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fintushel:2005:TDM

[Fin05] Ronald Fintushel. A tribute to Deane Montgomery. Notices of the American Mathematical Society, 52(3):348–349, March 2005. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Floyd:2006:HGT

[FK06] Juliet Floyd and Akihiro Kanamori. How Gödel transformed set theory. Notices of the American Mathematical Society, 53(4):419–427, April 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Friedlander:2004:OAL

[FLM^{+}04] Susan Friedlander, Peter Lax, Cathleen Morawetz, Louis Nirenberg, Gregory Seregin, Nina Ural’tseva, and Mark Vishik. Olga Alexandrovna Ladyzhenskaya (1922–2004). Notices of the American Mathematical Society, 51(11):1320– 1331, December 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fowler:2005:MAA

[Fow05] Kristine K. Fowler. The mathematics autodidact’s aid. Notices of the American Mathematical Society, 52(10):1203– 1207, November 2005. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Franzen:2006:PIG

[Fra06] Torkel Franzén. The popular impact of Gödel’s incompleteness theorem. Notices of the American Mathematical Society, 53(4):440–443, April 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Friedman:2000:FBP

[Fri00] Avner Friedman. Free boundary problems in science and technology. Notices of the American Mathematical Society, 47 (8):854–861, September 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Friedberg:2005:TMG

[Fri05] Solomon Friedberg. Teaching mathematics graduate students how to teach. Notices of the American Mathematical Society, 52(8):842–847, September 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Friberg:2008:RCB

[Fri08] Jöran Friberg. A remarkable collection of Babylonian mathematical texts. Notices of the American Mathematical Society, 55(9):1076–1086, October 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Friedlander:2003:MWF

[FRS03] Eric M. Friedlander, Michael Rapoport, and Andrei Suslin. The mathematical work of the 2002 Fields medalists. Notices of the American Mathematical Society, 50(2):212–217, February 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fokas:2005:GFT

[FS05] A. S. Fokas and L.-Y. Sung.

Generalized Fourier transforms, their nonlinearization and the imaging of the brain. Notices of the American Mathematical Society, 52(10):1178– 1192, November 2005. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gale:2007:LPS

[Gal07] David Gale. Linear programming and the simplex method. Notices of the American Mathematical Society, 54(3):364–369, March 2007. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Golomb:2002:CES

[GBC^{+}02] Solomon W. Golomb, Elwyn Berlekamp, Thomas M. Cover, Robert G. Gallager, James L. Massey, and Andrew J. Viterbi. Claude Elwood Shannon (1916– 2001). Notices of the American Mathematical Society, 49(1):8– 16, January 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/
notices/200201/fea-shannon.
pdf.

Gelfand:2009:WDC

[Gel09] Mikhail Gelfand. We do not choose mathematics as our profession, it chooses us: interview with Yuri Manin. Notices of the American Mathematical Society, 56(10):1268– 1274, November 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Translated from the Russian

http://www.polit.ru/science/2008/10/16/manin.html

by Mark Saul.

Grattan-Guinness:2000:SLH

[GG00] Ivor Grattan-Guinness. A sideways look at Hilbert’s twentythree problems of 1900. Notices of the American Mathematical Society, 47(7):752–757, August 2000. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Greuel:2007:MWJ

[GHW07] Gert-Martin Greuel, John E.

Hopcroft, and Margaret H. Wright. The mathematical work of Jon Kleinberg. Notices of the American Mathematical Society, 54(6):740–743, June/July 2007. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gindikin:2009:WLC

[Gin09] Simon Gindikin. What is … the complex dual to the real sphere? Notices of the American Mathematical Society, 56(8):942–943, September 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goel:2006:IMA

[Goe06] Sharad Goel. An invisible minority: Asian Americans in mathematics. Notices of the American Mathematical Society, 53(8):878–882, September 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goldman:2007:WLP

[Gol07] William M. Goldman. What is … a projective structure?

Notices of the American Mathematical Society, 54(1):30–33, January 2007. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Golden:2009:CCM

[Gol09a] Kenneth M. Golden. Climate change and the mathematics of transport in sea ice. Notices of the American Mathematical Society, 56(5):562–584, May 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Goldin:2009:DDB

[Gol09b] Rebecca Goldin. Doting on data. Book review of Super crunchers: why thinking-by-numbers is the new way to be smart. Notices of the American Mathematical Society, 56 (4):483–487, April 2009. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gompf:2005:WLL

[Gom05] Robert E. Gompf. What is … a Lefschetz pencil? Notices of the American Mathematical Society, 52(8):848–850, September 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gonthier:2008:FPF

[Gon08] Georges Gonthier. Formal proof — the four-color theorem. Notices of the American Mathematical Society, 55(11):1382– 1393, December 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gray:2003:MLW

[Gra03] Lawrence Gray. A mathematician looks at S. Wolfram’s new kind of science. Review of: A new kind of science [Wolfram Media, Inc., Champaign, IL, 2002; 1920418]. Notices of the American Mathematical Society, 50(2):200–211, February 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gratzer:2007:TPS

[Grä07] George Grätzer. Two problems that shaped a century of lattice theory. Notices of the American Mathematical Society, 54 (6):696–707, June/July 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gray:2008:CCT

[Gra08] Shirley B. Gray. A centennial celebration of two great scholars: Heiberg’s translation of the lost palimpsest of Archimedes — 1907. Heath’s publication on Euclid’s Elements — 1908. Notices of the American Mathematical Society, 55(7):776–783, August 2008. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gratzer:2009:WNLa

[Grä09a]  George Grätzer.  What is

new in LaTEX? I. Breaking free. Notices of the American Mathematical Society, 56(1):52– 54, January 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/
notices/200901/tx090100052p.
pdf.

Gratzer:2009:WNLb

[Grä09b] George Grätzer. What is new in LaTEX? II. TEX implementations, evolution or revolution? Notices of the American Mathematical Society, 56(5): 627–629, May 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/
notices/200905/rtx090500627p.
pdf.

Gratzer:2009:WNLc

[Grä09c]  George Grätzer.  What is

new in LaTEX? III. Formatting references. Notices of the American Mathematical Society, 56(8):954–956, September 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/
notices/200908/rtx090800954p.
pdf.

Grossberg:2000:LMB

[Gro00] Stephen Grossberg. Linking mind to brain: the mathematics of biological intelligence. Notices of the American Mathematical Society, 47(11):1361– 1372, December 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Grossman:2005:PRM

[Gro05] Jerrold W. Grossman. Patterns of research in mathematics. Notices of the American Mathematical Society, 52(1):35– 41, January 2005. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Grunbaum:2006:WSG

[Grü06] Branko Grünbaum. What symmetry groups are present in the Alhambra? Notices of the American Mathematical Society, 53(6):670–673, June/July 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gritzmann:2008:VK

[GS08] Peter Gritzmann and Bernd Sturmfels. Victor L. Klee 1925– 2007. Notices of the American Mathematical Society, 55 (4):467–473, April 2008. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With contributions and an editors’ epilogue.

Granville:2002:E

[GT02] Andrew Granville and Thomas J.

Tucker. It’s as easy as abc. Notices of the American Mathematical Society, 49(10):1224– 1231, November 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gunnells:2006:CCG

[Gun06] Paul E. Gunnells. Cells in Coxeter groups. Notices of the American Mathematical Society, 53(5):528–535, May 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gunning:2009:MPP

[Gun09] Robert C. Gunning. The Millennium Prize Problems [book review of MR2246251]. Notices of the American Mathematical Society, 56(11):1438–1439, December 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Guo:2009:WLR

[Guo09] Li Guo. What is … a Rota– Baxter algebra? Notices of the American Mathematical Society, 56(11):1436–1437, December 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hales:2000:CH

[Hal00] Thomas C. Hales. Cannonballs and honeycombs. Notices of the American Mathematical Society, 47(4):440– 449, April 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/
notices/200004/fea-hales.

pdf. In January 2003, the author was awarded the Chauvenet Prize of the Mathematical Association of America for this article.

Hales:2008:FP

[Hal08] Thomas C. Hales. Formal proof. Notices of the American Mathematical Society, 55(11):1370– 1380, December 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/
notices/200811/index.html;
 http://www.ams.org/notices/
200811/tx081101370p.pdf.

Hartshorne:2000:TGA

[Har00] Robin Hartshorne. Teaching geometry according to Euclid. Notices of the American Mathematical Society, 47(4):460–465, April 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Harris:2004:SFB

[Har04] Michael Harris. A sometimes funny book supposedly about infinity — a review of Everything and more [Atlas Books, W. W. Norton & Company, New York, 2003; MR2016410] by D. F. Wallace. Notices of the American Mathematical Society, 51(6):632–638, June/ July 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Harrison:2008:FPT

[Har08] John Harrison. Formal proof — theory and practice. Notices of the American Mathematical Society, 55(11):1395–1406, December 2008. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hayes:2001:LC

[Hay01] David R. Hayes. Leonard Carlitz (1907–1999). Notices of the American Mathematical Society, 48(11):1322–1324, December 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hefferon:2009:TF

[HB09] Jim Hefferon and Karl Berry. The TEX family in 2009. Notices of the American Mathematical Society, 56(3):348–354, March 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://www.

                                        ams.org/journals/notices/
                                        200903/tx090300348p.pdf.

Heinonen:2006:WLQ

[Hei06] Juha Heinonen. What is … a quasiconformal mapping? Notices of the American Mathematical Society, 53(11):1334– 1335, December 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hironaka:2009:WLL

[Hir09] Eriko Hironaka. What is … Lehmer’s number? Notices of the American Mathematical Society, 56(3):374–375, March 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hirzebruch:2009:CGT

[HK09] Friedrich E. P. Hirzebruch and Matthias Kreck. On the concept of genus in topology and complex analysis. Notices of the American Mathematical Society, 56(6):713–719, June/July 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Herzberg:2007:SSC

[HM07] Agnes M. Herzberg and M. Ram Murty. Sudoku squares and chromatic polynomials. Notices of the American Mathematical Society, 54(6):708–717, June/ July 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hobbs:2004:WT

[HO04] Arthur M. Hobbs and James G.

Oxley. William T. Tutte (1917– 2002). Notices of the American Mathematical Society, 51 (3):320–330, March 2004. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hodges:2001:WDY

[Hod01] Wilfrid Hodges. What do you want from your publisher? An annotated checklist for mathematical authors. Notices of the American Mathematical Society, 48(10):1176–1181, November 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hodges:2006:ETB

[Hod06] Andrew Hodges. The Essential Turing — a book review. Notices of the American Mathematical Society, 53(10):1190– 1199, November 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

                                        URL http://www.ams.org/
                                        notices/200610/rev-hodges.
                                        pdf.

Horvath:2009:EMC

[Hor09] John Horváth. Encounters with Mischa Cotlar. Notices of the American Mathematical Society, 56(5):616–620, May 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Hardin:2004:DMM

[HS04] D. P. Hardin and E. B. Saff. Discretizing manifolds via minimum energy points. Notices of the American Mathematical Society, 51(10):1186–1194, November 2004. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Huckleberry:2009:HGM

[Huc09]  Alan Huckleberry.  Hans

Grauert: Mathematiker pur. Notices of the American Mathematical Society, 56(1):38–41, January 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Isaev:2000:IDM

[IK00] Alexander V. Isaev and Steven G. Krantz. Invariant distances and metrics in complex analysis. Notices of the American Mathematical Society, 47(5):546–553, May 2000. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Illusie:2004:WLT

[Ill04] Luc Illusie. What is … a topos? Notices of the American Mathematical Society, 51 (9):1060–1061, October 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Iosevich:2001:CCF

[Ios01] Alex Iosevich. Curvature, combinatorics, and the Fourier transform. Notices of the American Mathematical Society, 48 (6):577–583, June/July 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Iserles:2002:EGT

[Ise02] Arieh Iserles. Expansions that grow on trees. Notices of the American Mathematical Society, 49(4):430–440, April 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Its:2003:RHP

[Its03]  Alexander R. Its.  The

Riemann–Hilbert problem and integrable systems. Notices of the American Mathematical Society, 50(11):1389–1400, December 2003. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2000:EWR

[Jac00a] Allyn Jackson. Edward Witten receives Nemmers Prize. Notices of the American Mathematical Society, 47(6):666, June/ July 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2000:MBT

[Jac00b] Allyn Jackson. Mathematics in Barcelona: time past, time future. Notices of the American Mathematical Society, 47 (5):554–560, May 2000. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2000:NNI

[Jac00c] Allyn Jackson. Newest NSF institute opens its doors. Notices of the American Mathematical Society, 47(11):1412–1414, December 2000. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2000:OYT

[Jac00d] Allyn Jackson. Oberwolfach, yesterday and today. Notices of the American Mathematical Society, 47(7):758–765, August 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2000:RMR

[Jac00e] Allyn Jackson. Robert Moses receives Heinz Award. Notices of the American Mathematical Society, 47(8):884, September 2000. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2000:SRF

[Jac00f] Allyn Jackson. The slow revolution of the free electronic journal. Notices of the American Mathematical Society, 47 (9):1053–1059, October 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2001:IAR

[Jac01a] Allyn Jackson. Interview with Arnold Ross. Notices of the American Mathematical Society, 48(7):691–697, August 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2001:IRB

[Jac01b] Allyn Jackson. Interview with Raoul Bott. Notices of the American Mathematical Society, 48(4):374–382, April 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2001:PVI

[Jac01c] Allyn Jackson. Presidential views: interview with Hyman Bass. Notices of the American Mathematical Society, 48 (3):312–315, March 2001. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2002:ISM

[Jac02a] Allyn Jackson. The IAS School of Mathematics. Notices of the American Mathematical Society, 49(8):896–904, September 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2002:ILN

[Jac02b] Allyn Jackson. Interview with Louis Nirenberg. Notices of the American Mathematical Society, 49(4):441–449, April 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2002:WBM

[Jac02c] Allyn Jackson. The world of blind mathematicians. Notices of the American Mathematical Society, 49(10):1246–1251, November 2002. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2003:MCT

[Jac03a] Allyn Jackson. MSRI celebrates its twentieth birthday. Notices of the American Mathematical Society, 50(3):373–375, March 2003. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2003:PVIb

[Jac03b] Allyn Jackson. Presidential views: interview with David Eisenbud. Notices of the American Mathematical Society, 50 (3):370–372, March 2003. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2003:PVIa

[Jac03c] Allyn Jackson. Presidential views: interview with Hyman Bass. Notices of the American Mathematical Society, 50(2):232–234, February 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2003:STR

[Jac03d] Allyn Jackson. Sato and Tate receive 2002–2003 Wolf Prize. Notices of the American Mathematical Society, 50(5):569–570, May 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2003:SRA

[Jac03e] Allyn Jackson. Serre receives Abel Prize. Notices of the American Mathematical Society, 50 (6):693, June/July 2003. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2004:CADa

[Jac04a] Allyn Jackson. Comme appelé du néant — as if summoned from the void: the life of Alexandre Grothendieck. Notices of the American Mathematical Society, 51(9):1038– 1056, October 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2004:CADb

[Jac04b] Allyn Jackson. Comme appelé du néant — as if summoned from the void: the life of Alexandre Grothendieck. II. Notices of the American Mathematical Society, 51(10):1196– 1212, November 2004. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2004:GRN

[Jac04c] Allyn Jackson. Gromov receives Nemmers Prize. Notices of the American Mathematical Society, 51(7):787–788, August 2004. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2004:HHC

[Jac04d] Allyn Jackson. Happy 100th Henri Cartan! Notices of the American Mathematical Society, 51(7):771, August 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2004:IJK

[Jac04e] Allyn Jackson. Interview with Joseph Keller. Notices of the American Mathematical Society, 51(7):751–760, August 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2004:MP

[Jac04f] Allyn Jackson. Mathematics people. Notices of the American Mathematical Society, 51 (3):343–344, March 2004. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2004:OCS

[Jac04g] Allyn Jackson. Oberwolfach celebrates its sixtieth anniversary. Notices of the American Mathematical Society, 51 (9):1064–1066, October 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2004:RIJ

[Jac04h] Allyn Jackson. RIMS, an institute for Japan and the world. Notices of the American Mathematical Society, 51(2):194–200, February 2004. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2004:VRN

[Jac04i] Allyn Jackson. Voiculescu receives NAS award in mathematics. Notices of the American Mathematical Society, 51 (5):547, May 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2005:DKI

[Jac05a] Allyn Jackson. A different kind of institute: the American Institute of Mathematics. Notices of the American Mathematical Society, 52(11):1330–1337, December 2005. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2005:MC

[Jac05b]  Allyn Jackson.  In memo-

riam: S.-S. Chern. Notices of the American Mathematical Society, 52(6):623–625, June/ July 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2005:IHH

[Jac05c] Allyn Jackson. Interview with Heisuke Hironaka. Notices of the American Mathematical Society, 52(9):1010–1019, October 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2005:PVI

[Jac05d] Allyn Jackson. Presidential views: Interview with James Arthur. Notices of the American Mathematical Society, 52 (3):350–352, March 2005. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2005:WRS

[Jac05e] Allyn Jackson. Wiles receives 2005 Shaw Prize. Notices of the American Mathematical Society, 52(8):869, September 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2006:GP

[Jac06a] Allyn Jackson. 2006 Gauss Prize. Notices of the American Mathematical Society, 53(10): 1224–1225, November 2006.

CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2006:CNM

[Jac06b] Allyn Jackson. Conjectures no more? Consensus forming on the proof of the Poincaré and geometrization conjectures. Notices of the American Mathematical Society, 53(8):897–901, September 2006. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2006:DNR

[Jac06c] Allyn Jackson. Donaldson and Narasimhan receive 2006 King Faisal Prize. Notices of the American Mathematical Society, 53(3):347, March 2006. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2006:HGS

[Jac06d] Allyn Jackson. In high gear:

Spanish mathematics looks to the future — and to ICM2006. Notices of the American Mathematical Society, 53(2):218–222, February 2006. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2006:SRN

[Jac06e] Allyn Jackson. Sullivan receives 2004 National Medal of Science. Notices of the American Mathematical Society, 53 (3):346, March 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2007:BLF

[Jac07a] Allyn Jackson. Bers library finds a home — and resonance — at Charles University. Notices of the American Mathematical Society, 54(7):856–858, August 2007. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2007:ERN

[Jac07b]  Allyn Jackson.  Efron re-

ceives National Medal of Science. Notices of the American Mathematical Society, 54 (8):999, September 2007. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2008:GI

[Jac08a] Allyn Jackson. Grothendieck at 80, IHES at 50. Notices of the American Mathematical Society, 55(8):962–963, September 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2008:IMD

[Jac08b] Allyn Jackson. Interview with Martin Davis. Notices of the American Mathematical Society, 55(5):560–571, May 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2009:DMN

[Jac09a] Allyn Jackson. Donald McClure named AMS executive director.

Notices of the American Mathematical Society, 56(3):386–387, March 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2009:JER

[Jac09b] Allyn Jackson. John Ewing retires from the AMS. Notices of the American Mathematical Society, 56(1):48– 51, January 2009. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL https://www.
ams.org/journals/notices/
200901/200901-full-issue.

pdf. With contributions by Jonathan Borwein and Jane Kister. Borwein’s is a sidebar on page 49, with the title Some Reflections on John Ewing.

Jackson:2009:PVI

[Jac09c] Allyn Jackson. Presidential views: interview with James Glimm. Notices of the American Mathematical Society, 56(2): 247–249, February 2009. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2009:TI

[Jac09d] Allyn Jackson. A trio of institutes. Notices of the American Mathematical Society, 56 (11):1426–1434, December 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jaffe:2006:MGC

[Jaf06] Arthur M. Jaffe. The millennium grand challenge in mathematics. Notices of the American Mathematical Society, 53 (6):652–660, June/July 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jech:2008:WLF

[Jec08] Thomas Jech. What is … forcing? Notices of the American Mathematical Society, 55 (6):692–693, June/July 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jorgenson:2006:SL

[JK06] Jay Jorgenson and Steve G. Krantz. Serge Lang, 1927–2005. Notices of the American Mathematical Society, 53(5):536–553, May 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jorgenson:2007:MCS

[JK07] Jay Jorgenson and Steven G. Krantz. The mathematical contributions of Serge Lang. Notices of the American Mathematical Society, 54(4):476–497, April 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jager:2003:OAO

[JLM03] Willi Jäger, Peter Lax, and Cathleen Synge Morawetz.

Olga Arsen’evna Oleĭnik (1925– 2001). Notices of the American Mathematical Society, 50(2): 220–223, February 2003. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Johnson:2009:LBR

[Joh09] Jesse Johnson. Out of the labyrinth [book review of MR2279945]. Notices of the American Mathematical Society, 56(6):723–725, June/July 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jorgensen:2003:MFA

[Jor03] Palle E. T. Jorgensen. Matrix factorizations, algorithms, wavelets. Notices of the American Mathematical Society, 50 (8):880–894, September 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:2003:AR

[JS03] Allyn Jackson and Daniel Shapiro. Arnold Ross (1906– 2002). Notices of the American Mathematical Society, 50 (6):660–665, June/July 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With contributions by: Prakash Bambah, Thomas Banchoff, Felix Browder, David Pollack, Peter Roquette, Karl Rubin, Paul J. Sally Jr., James Schultz, Alice Silverberg, Glenn Stevens, Bert K. Waits, Max Warshauer and Gloria Woods.

Joyner:2007:OSM

[JS07] David Joyner and William Stein. Open source mathematical software. Notices of the American Mathematical Society, 54(10):1279, November 2007. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

                                        URL http://www.ams.org/
                                        notices/200710/tx071001279p.
                                        pdf;   http://www.sagemath.
                                        org/.

Johnson:2006:SMA

[JW06] G. W. Johnson and Mark E. Walker. Sir Michael Atiyah’s Einstein lecture: “The nature of space”. Notices of the American Mathematical Society, 53(6):674–678, June/ July 2006. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

                                        URL http://www.ams.org/
                                        notices/200606/comm-walker.
                                        pdf.

Kadison:2008:IKR

[Kad08] Richard V. Kadison. Irving Kaplansky’s role in mid-twentieth century functional analysis. Notices of the American Mathematical Society, 55(2):216–225, February 2008. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Karp:2002:MCG

[Kar02] Richard M. Karp. Mathematical challenges from genomics and molecular biology. Notices of the American Mathematical Society, 49(5):544–553, May 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kesten:2006:WLP

[Kes06] Harry Kesten. What is … percolation? Notices of the American Mathematical Society, 53 (5):572–573, May 2006. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kohn:2004:DCS

[KGG^{+}04] Joseph J. Kohn, Phillip A. Griffiths, Hubert Goldschmidt, Enrico Bombieri, Bohous Cenkl, Paul Garabedian, and Louis Nirenberg. Donald C. Spencer (1912–2001). Notices of the American Mathematical Society, 51(1):17–29, January 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Khesin:2005:TFD

[Khe05] Boris Khesin. Topological fluid dynamics. Notices of the American Mathematical Society, 52 (1):9–19, January 2005. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kirby:2002:CAM

[Kir02] Rob Kirby. Conversations about mathematics. Notices of the American Mathematical Society, 49(3):333–335, March 2002.

CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kisil:2007:SG

[Kis07a] Vladimir V. Kisil. Starting with the group SL_{2}(R). Notices of the American Mathematical Society, 54(11):1458–1465, December 2007. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kisin:2007:WLG

[Kis07b] Mark Kisin. What is … a Galois representation? Notices of the American Mathematical Society, 54(6):718–719, June/July 2007. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Khavinson:2008:FTA

[KN08] Dmitry Khavinson and Genevra Neumann. From the fundamental theorem of algebra to astrophysics: A “Harmonious” path. Notices of the American Mathematical Society, 55 (6):666–675, June/July 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

                                        tronic). URL http://www.
                                        ams.org/journals/notices/
                                        200806/noti-jun08-cov1.jpg.

See follow-on article [Pet10].

Knuth:2002:AQA

[Knu02] Donald Knuth. All questions answered. Notices of the American Mathematical Society, 49 (3):318–324, March 2002. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kenyon:2005:WLD

[KO05] Richard Kenyon and Andrei Okounkov. What is … a dimer? Notices of the American Mathematical Society, 52(3):342–343, March 2005. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Koblitz:2007:URB

[Kob07] Neal Koblitz. The uneasy relationship between mathematics and cryptography. Notices of the American Mathematical Society, 54(8):972–979, September 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kollar:2006:TM

[Kol06] János Kollár. Teruhisa Matsusaka (1926–2006). Notices of the American Mathematical Society, 53(7):766–768, August 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kollar:2007:WLM

[Kol07] János Kollár. What is … a minimal model? Notices of the American Mathematical Society, 54(3):370–371, March 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kopell:2000:WGR

[Kop00] Nancy Kopell. We got rhythm:

dynamical systems of the nervous system. Notices of the American Mathematical Society, 47(1):6–16, January 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kotschick:2004:WLQ

[Kot04] D. Kotschick. What is … a quasi-morphism? Notices of the American Mathematical Society, 51(2):208–209, February 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Krantz:2007:HWY

[Kra07] Steven G. Krantz. How to write your first paper. Notices of the American Mathematical Society, 54(11):1507–1511, December 2007. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Krieger:2004:SWM

[Kri04] Martin H. Krieger. Some of what mathematicians do. Notices of the American Mathematical Society, 51(10):1226– 1230, November 2004. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Krieger:2005:LAW

[Kri05] Martin H. Krieger. A 1940 letter of André Weil on analogy in mathematics. Notices of the American Mathematical Society, 52(3):334–341, March 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Excerpted from Doing mathematics [World Scientific Publishing Co., Inc., River Edge, NJ, 2003; MR1961400].

Kister:2004:RGB

[KS04] Jane E. Kister and Donald R.

Sherbert. Robert G. Bartle (1927–2003). Notices of the American Mathematical Society, 51(2):239–240, February 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kuo:2005:LCD

[KS05] Frances Y. Kuo and Ian H.

Sloan. Lifting the curse of dimensionality. Notices of the American Mathematical Society, 52(11):1320–1329, December 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kosmann-Schwarzbach:2009:TAL

[KS09] Yvette Kosmann-Schwarzbach.

Tribute to André Lichnerowicz (1915–1998). Notices of the American Mathematical Society, 56(2):244–246, February 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Knutson:2001:HSH

[KT01] Allen Knutson and Terence Tao.

Honeycombs and sums of Hermitian matrices. Notices of the American Mathematical Society, 48(2):175–186, February 2001. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Klink:2009:ITT

[KTT09] William H. Klink and Tuong Ton-That. Invariant theory of tensor product decompositions of U(N) and generalized Casimir operators. Notices of the American Mathematical Society, 56(8):931–940, September 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kurshan:2000:PV

[Kur00] Robert P. Kurshan. Program verification. Notices of the American Mathematical Society, 47(5):534–545, May 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Labourie:2008:WLC

[Lab08] François Labourie. What is … a cross ratio? Notices of the American Mathematical Society, 55(10):1234–1235, November 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Landau:2000:CST

[Lan00a] Susan Landau. Communications security for the Twentyfirst Century: The Advanced Encryption Standard. Notices of the American Mathematical Society, 47(4):450– 459, April 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/
notices/200004/fea-landau.
pdf.

Landau:2000:STT

[Lan00b]  Susan Landau.  Standing

the test of time: the Data Encryption Standard. Notices of the American Mathematical Society, 47(3):341– 349, March 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/
notices/200003/fea-landau.
pdf.

Landsburg:2004:QGT

[Lan04] Steven E. Landsburg. Quantum game theory. Notices of the American Mathematical Society, 51(4):394–399, April 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lang:2005:ANP

[Lan05] Serge Lang. On the AMS Notices publication of Krieger’s translation of Weil’s 1940 letter [MR2125268]. Notices of the American Mathematical Society, 52(6):612–622, June/July 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Landau:2006:FMH

[Lan06] Susan Landau. Find me a hash. Notices of the American Mathematical Society, 53(3): 330–332, March 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

                                        URL http://www.ams.org/
                                        notices/200603/fea-landau.
                                        pdf.

Lax:2003:MS

[Lax03] Peter D. Lax. Max Shiffman (1914–2000). Notices of the American Mathematical Society, 50(11):1401, December 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lenstra:2002:SPE

[Len02] H. W. Lenstra, Jr. Solving the Pell equation. Notices of the American Mathematical Society, 49(2):182–192, February 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Leonard:2008:AAY

[Leo08] Kathryn Leonard. Adventures in academic year undergraduate research. Notices of the American Mathematical Society, 55 (11):1422–1426, December 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lickorish:2009:WLS

[Lic09] W. B. Raymond Lickorish. What is … a skein module? Notices of the American Mathematical Society, 56(2):240–242, February 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lipscomb:2009:QUS

[Lip09] Stephen Leon Lipscomb. The quest for universal spaces in dimension theory. Notices of the American Mathematical Society, 56(11):1418–1424, December 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lax:2001:JLL

[LMT01] Peter D. Lax, Enrico Magenes, and Roger Temam. Jacques-Louis Lions (1928–2001). Notices of the American Mathematical Society, 48(11):1315– 1321, December 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lubotzky:2004:MBW

[Lub04] Alexander Lubotzky. The mathematics of Boris Weisfeiler. Notices of the American Mathematical Society, 51(1):31–32, January 2004. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lubotzky:2005:WLP

[Lub05] Alex Lubotzky. What is … property (τ)? Notices of the American Mathematical Society, 52(6):626–627, June/July 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lurie:2008:WLC

[Lur08] Jacob Lurie. What is … an ∞- category? Notices of the American Mathematical Society, 55 (8):949–950, September 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lyubich:2000:QFQ

[Lyu00] Mikhail Lyubich. The quadratic family as a qualitatively solvable model of chaos. Notices of the American Mathematical Society, 47(9):1042–1052, October 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Majid:2006:WLQ

[Maj06] Shahn Majid. What is … a quantum group? Notices of the American Mathematical Society, 53(1):30–31, January 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Marateck:2006:DGP

[Mar06] Samuel L. Marateck. The differential geometry and physical basis for the applications of Feynman diagrams. Notices of the American Mathematical Society, 53(7):744– 752, August 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/
notices/200607/fea-marateck.
pdf.

Mawhin:2005:HPL

[Maw05] Jean Mawhin. Henri Poincaré. A life in the service of science. Notices of the American Mathematical Society, 52 (9):1036–1044, October 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Maxwell:2008:WJL

[Max08] James W. Maxwell. William J. LeVeque (1923–2007). Notices of the American Mathematical Society, 55(10):1261– 1262, November 2008. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mazur:2000:QAP

[Maz00] Barry Mazur. Questions about powers of numbers. Notices of the American Mathematical Society, 47(2):195–202, February 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mazur:2004:WLM

[Maz04] Barry Mazur. What is … a motive? Notices of the American Mathematical Society, 51(10):1214–1216, November 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

McColm:2007:MME

[McC07] Greg McColm. A metaphor for mathematics education. Notices of the American Mathematical Society, 54(4):499–502, April 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

McGeoch:2001:EAA

[McG01] Catherine C. McGeoch. Experimental analysis of algorithms. Notices of the American Mathematical Society, 48(3):304–311, March 2001. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mezzadri:2007:HGR

[Mez07] Francesco Mezzadri. How to generate random matrices from the classical compact groups. Notices of the American Mathematical Society, 54(5):592–604, May 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mikhalkin:2007:WLT

[Mik07] Grigory Mikhalkin. What is … a tropical curve? Notices of the American Mathematical Society, 54(4):511–513, April 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Milnor:2003:TPC

[Mil03] John Milnor. Towards the Poincaré conjecture and the classification of 3-manifolds. Notices of the American Mathematical Society, 50(10):1226– 1233, November 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Miller:2008:WLT

[Mil08a] Ezra Miller. What is … a toric variety? Notices of the American Mathematical Society, 55 (5):586–587, May 2008. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Miller:2008:CFG

[Mil08b] Russell Miller. Computable fields and Galois theory. Notices of the American Mathematical Society, 55(7):798–807, August 2008. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Miner:2005:IMM

[Min05] Robert Miner. The importance of MathML to mathematics communication. Notices of the American Mathematical Society, 52(5):532–538, May 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mather:2000:JKM

[MMNR00] John N. Mather, Henry P.

McKean, Louis Nirenberg, and Paul H. Rabinowitz. Jürgen K. Moser (1928–1999). Notices of the American Mathematical Society, 47(11):1392–1405, December 2000. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mohar:2006:WLG

[Moh06] Bojan Mohar. What is … a graph minor. Notices of the American Mathematical Society, 53(3):338–339, March 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Moll:2002:EIP

[Mol02] Victor H. Moll. The evaluation of integrals: a personal story. Notices of the American Mathematical Society, 49(3):311–317, March 2002. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Montgomery:2001:NST

[Mon01] Richard Montgomery. A new solution to the three-body problem. Notices of the American Mathematical Society, 48 (5):471–481, May 2001. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Moore:2005:WLB

[Moo05] Gregory W. Moore. What is … a brane? Notices of the American Mathematical Society, 52(2):214–215, February 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Moore:2009:EZA

[Moo09]  Gregory H. Moore.  Ernst

Zermelo: an approach to his life and work [book review of MR2301067]. Notices of the American Mathematical Society, 56(7):823–827, August 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Morgan:2005:BRB

[Mor05a] Frank Morgan. Book review: Kepler’s Conjecture and Hales’ Proof. Notices of the American Mathematical Society, 52(1):44– 47, January 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/
notices/200501/rev-morgan.

pdf. Review of [Szp03].

Morgan:2005:MD

[Mor05b] Frank Morgan. Manifolds with density. Notices of the American Mathematical Society, 52 (8):853–858, September 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Munoz:2007:ITF

[MP07] Vicente Muñoz and Ulf Persson.

Interviews with three Fields medalists. Notices of the American Mathematical Society, 54 (3):405–410, March 2007. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Maslen:2001:CTF

[MR01] David K. Maslen and Daniel N.

Rockmore. The Cooley–Tukey FFT and group theory. Notices of the American Mathematical Society, 48(10):1151– 1160, November 2001. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mackey:2005:MBP

[MS05] Michael C. Mackey and Moisés Santillán. Mathematics, biology, and physics: interactions and interdependence. Notices of the American Mathematical Society, 52(8):832–840, September 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mycielski:2006:SAS

[Myc06] Jan Mycielski. A system of axioms of set theory for the rationalists. Notices of the American Mathematical Society, 53(2):206–213, February 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Neuhauser:2001:MCS

[Neu01] Claudia Neuhauser. Mathematical challenges in spatial ecology. Notices of the American Mathematical Society, 48(11):1304– 1314, December 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Neyman:2006:AAN

[Ney06] Abraham Neyman. Aumann awarded Nobel Prize. Notices of the American Mathematical Society, 53(1):44–46, January 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Nowak:2008:WLP

[NY08] Piotr Nowak and Guoliang Yu. What is … property A? Notices of the American Mathematical Society, 55(4):474–475, April 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ochanine:2009:WIE

[Och09] Serge Ochanine. What is … an elliptic genus? Notices of the American Mathematical Society, 56(6):720–721, June/ July 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ono:2006:HGK

[Ono06] Ken Ono. Honoring a gift from Kumbakonam. Notices of the American Mathematical Society, 53(6):640–651, June/ July 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Osserman:2005:MH

[Oss05] Robert Osserman. Mathematics of the heavens. Notices of the American Mathematical Society, 52(4):417–424, April 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ovsienko:2009:WLS

[OT09] Valentin Ovsienko and Sergei Tabachnikov. What is … the Schwarzian derivative? Notices of the American Mathematical Society, 56(1):34–36, January 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pitman:2005:MSP

[PA05] Jim Pitman and David Aldous. The Mathematics Survey Project. Notices of the American Mathematical Society, 52(11):1357–1360, December 2005. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/
notices/200511/comm-pitman.
pdf.

Porter:2005:GCN

[PC05] Mason A. Porter and Predrag Cvitanović. Ground control to Niels Bohr: exploring outer space with atomic physics. Notices of the American Mathematical Society, 52(9):1020– 1025, October 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Perlick:2000:ROF

[Per00] Volker Perlick. Ray Optics, Fermat’s Principle, and Applications to General Relativity, volume 61 of Lecture notes in physics: New series m, Monographs; Physics and astronomy online library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-66898-5. x + 220 pp. LCCN QC389 .P37 2000.

Persson:2005:IAS

[Per05] Ulf Persson. Interview with Arild Stubhaug. Notices of the American Mathematical Society, 52(9):1046–1051, October 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Peters:2009:WPM

[Pet09]  Klaus Peters.  Why pub-

lish mathematics? Notices of the American Mathematical Society, 56(7):819– 822, August 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Reprinted from Publikationsstrategien einer Disziplin Harrassowitz Verlag, Wiesbaden, 2008.

Petters:2010:GAL

[Pet10]  Arlie O. Petters.  Grav-

ity’s action on light. Notices of the American Mathematical Society, 57(11):1392– 1409, December 2010. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL http://www.
ams.org/journals/notices/
201011/201011-full-issue.

pdf. See preceding article [KN08], and book-length treatment [Per00].

Pournaki:2009:WLS

[PFTY09] M. R. Pournaki, S. A. Seyed Fakhari, M. Tousi, and S. Yassemi. What is … Stanley depth? Notices of the American Mathematical Society, 56(9):1106– 1108, October 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pillay:2000:MT

[Pil00] Anand Pillay. Model theory.

Notices of the American Mathematical Society, 47(11):1373– 1381, December 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Porter:2006:MB

[PL06] Mason A. Porter and Steven Lansel. Mushroom billiards. Notices of the American Mathematical Society, 53(3):334–337, March 2006. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Poenaru:2007:WLI

[Poé07] Valentin Poénaru. What is … an infinite swindle? Notices of the American Mathematical Society, 54(5):619–622, May 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Poenaru:2008:MS

[Poé08] Valentin Poénaru. Memories of Shourik. Notices of the American Mathematical Society, 55 (8):964–965, September 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Porter:2009:CN

[POM09] Mason A. Porter, Jukka-Pekka Onnela, and Peter J. Mucha. Communities in networks. Notices of the American Mathematical Society, 56(9):1082– 1097, October 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Poonen:2008:UNT

[Poo08] Bjorn Poonen. Undecidability in number theory. Notices of the American Mathematical Society, 55(3):344–350, March 2008. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Protter:2007:WKI

[Pro07] Philip Protter. The work of Kyosi Itô. Notices of the American Mathematical Society, 54 (6):744–745, June/July 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Poggio:2003:MLD

[PS03] Tomaso Poggio and Steve Smale. The mathematics of learning: dealing with data. Notices of the American Mathematical Society, 50(5):537–544, May 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Pfender:2004:KNS

[PZ04] Florian Pfender and Günter M.

Ziegler. Kissing numbers, sphere packings, and some unexpected proofs. Notices of the American Mathematical Society, 51 (8):873–883, September 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Quarteroni:2009:MMS

[Qua09] Alfio Quarteroni. Mathematical models in science and engineering. Notices of the American Mathematical Society, 56(1):10– 19, January 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Reed:2004:WMB

[Ree04] Michael C. Reed. Why is mathematical biology so hard? Notices of the American Mathematical Society, 51(3):338–342, March 2004. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Reitberger:2002:LV

[Rei02] Heinrich Reitberger. Leopold Vietoris (1891–2002). Notices of the American Mathematical Society, 49(10):1232–1236, November 2002. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Riehm:2002:EHF

[Rie02] Carl Riehm. The early history of the Fields Medal. Notices of the American Mathematical Society, 49(7):778–782, August 2002. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rajapakse:2008:HSD

[RMM08] Indika Rajapakse, Lindsey Muir, and Paul Martin. Hardy’s “small” discovery remembered. Notices of the American Mathematical Society, 55(3):384–386, March 2008. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Roe:2006:WLC

[Roe06] John Roe. What is … a coarse space? Notices of the American Mathematical Society, 53 (6):668–669, June/July 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rogawski:2000:NRL

[Rog00] Jonathan Rogawski. The nonabelian reciprocity law for local fields. Notices of the American Mathematical Society, 47(1):35– 41, January 2000. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rosenau:2005:WLC

[Ros05] Philip Rosenau. What is … a compacton? Notices of the American Mathematical Society, 52(7):738–739, August 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rowe:2001:LBB

[Row01] David E. Rowe. Looking back on a bestseller: D. Struik’s A concise history of mathematics [Dover, New York, 1987; MR0919604 (88i:01004)]. Notices of the American Mathematical Society, 48(6):590–592, June/July 2001. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Raussen:2004:IJP

[RS04] Martin Raussen and Christian Skau. Interview with Jean-Pierre Serre. Notices of the American Mathematical Society, 51(2):210–214, February 2004. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Raussen:2005:IMA

[RS05] Martin Raussen and Christian Skau. Interview with Michael Atiyah and Isadore Singer. Notices of the American Mathematical Society, 52(2):225–233, February 2005. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Raussen:2006:IPD

[RS06] Martin Raussen and Christian Skau. Interview with Peter D. Lax. Notices of the American Mathematical Society, 53(2):223–229, February 2006. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

                                        URL http://www.ams.org/
                                        notices/200602/comm-lax.pdf.

Raussen:2007:IAP

[RS07] Martin Raussen and Christian Skau. Interview with Abel Prize recipient Lennart Carleson. Notices of the American Mathematical Society, 54(2):223–229, February 2007. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Raussen:2008:ISV

[RS08] Martin Raussen and Christian Skau. Interview with Srinivasa Varadhan. Notices of the American Mathematical Society, 55(2):238–246, February 2008. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic).

Raussen:2009:IJG

[RS09] Martin Raussen and Christian Skau. Interview with John G. Thompson and Jacques Tits. Notices of the American Mathematical Society, 56(4):471–478, April 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ruskey:2006:SSS

[RSW06] Frank Ruskey, Carla D. Savage, and Stan Wagon. The search for simple symmetric Venn diagrams. Notices of the American Mathematical Society, 53 (11):1304–1312, December 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rudnick:2008:WLQ

[Rud08] Ze’ev Rudnick. What is … quantum chaos? Notices of the American Mathematical Society, 55(1):32–34, January 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ruelle:2002:DZF

[Rue02] David Ruelle. Dynamical zeta functions and transfer operators. Notices of the American Mathematical Society, 49 (8):887–895, September 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ruelle:2006:WLS

[Rue06] David Ruelle. What is … a strange attractor? Notices of the American Mathematical Society, 53(7):764–765, August 2006. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rutter:2002:MML

[Rut02] Sara Rutter. Mathematicians and the mathematics library: a librarian’s perspective. Notices of the American Mathematical Society, 49(9):1078–1081, October 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Saari:2008:MV

[Saa08] Donald G. Saari. Mathematics and voting. Notices of the American Mathematical Society, 55(4):448–455, April 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sabloff:2009:WLL

[Sab09] Joshua M. Sabloff. What is … a Legendrian knot? Notices of the American Mathematical Society, 56(10):1282– 1284, November 2009. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sarnak:2000:RP

[Sar00] Peter Sarnak. Ralph Phillips (1913–1998). Notices of the American Mathematical Society, 47(5):561–563, May 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sarnak:2004:WLE

[Sar04] Peter Sarnak. What is … an expander? Notices of the American Mathematical Society, 51 (7):762–763, August 2004. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sastri:2009:MCS

[Sas09] Chelluri C. A. Sastri. Mathematics and common sense: a case of creative tension [book review of MR2273791]. Notices of the American Mathematical Society, 56(6):726–728, June/July 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Saul:2000:ACL

[Sau00]  Mark Saul.  Anneli Cahn

Lax (1922–1999). Notices of the American Mathematical Society, 47(7):766–769, August 2000. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Saul:2003:MSP

[Sau03] Mark Saul. Mathematics in a small place: notes on the mathematics of Romania and Bulgaria. Notices of the American Mathematical Society, 50 (5):561–565, May 2003. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Saloff-Coste:2001:PGR

[SC01] Laurent Saloff-Coste. Probability on groups: random walks and invariant diffusions. Notices of the American Mathematical Society, 48(9):968–977, October 2001. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Schapira:2007:MSV

[Sch07] Pierre Schapira. Mikio Sato, a visionary of mathematics. Notices of the American Mathematical Society, 54(2):243–245, February 2007. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Scharlau:2008:WAG

[Sch08] Winfried Scharlau. Who is Alexander Grothendieck? Notices of the American Mathematical Society, 55(8):930–941, September 2008. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Translated from the 2007 German original by D. Kotschick with the assistance of Allyn Jackson.

Schochet:2009:DDI

[Sch09] Claude Schochet. The Dixmier– Douady invariant for dummies. Notices of the American Mathematical Society, 56(7):809–816, August 2009. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Semmes:2003:IAM

[Sem03a] Stephen Semmes. An introduction to analysis on metric spaces. Notices of the American Mathematical Society, 50 (4):438–443, April 2003. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Semmes:2003:IHG

[Sem03b] Stephen Semmes. An introduction to Heisenberg groups in analysis and geometry. Notices of the American Mathematical Society, 50(6):640–646, June/July 2003. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Senechal:2006:WLQ

[Sen06] Marjorie Senechal. What is … a quasicrystal? Notices of the American Mathematical Society, 53(8):886–887, September 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Serfati:2008:NGD

[Ser08] Michel Serfati. A note on the Geometry and Descartes’s mathematical work. Notices of the American Mathematical Society, 55(1):50–53, January 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Shalev:2001:AGT

[Sha01] Aner Shalev. Asymptotic group theory. Notices of the American Mathematical Society, 48 (4):383–389, April 2001. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Shallit:2005:BRM

[Sha05]  Jeffrey Shallit.  Book re-

view: Mathematics by Experiment, Jonathan Borwein and David Bailey, A K Peters, 2003, 288 pages, $45.00, ISBN 1-56881-211-6. Experimentation in Mathematics, Jonathan Borwein, David Bailey, and Roland Girgensohn, A K Peters, 2004, 357 pages, $49.00, ISBN 1-56881-136-5. Notices of the American Mathematical Society, 52(8):863– 865, September 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL http://www.ams.org/
notices/200508/rev-shallit.
pdf.

Shallit:2009:NBR

[Sha09] Jeffrey Shallit. The numerati [book review; Houghton Mifflin, Geneva, IL, 2008]. Notices of the American Mathematical Society, 56(9):1109–1111, October 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Shub:2005:WLH

[Shu05] Michael Shub. What is … a horseshoe? Notices of the American Mathematical Society, 52(5):516–517, May 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sigmund:2006:PE

[Sig06] Karl Sigmund. Pictures at an exhibition. Notices of the American Mathematical Society, 53 (4):428–432, April 2006. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Silver:2008:LPJ

[Sil08] Daniel S. Silver. The last poem of James Clerk Maxwell. Notices of the American Mathematical Society, 55(10):1266– 1270, November 2008. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sinai:2004:WLB

[Sin04] Yakov Sinai. What is … a billiard? Notices of the American Mathematical Society, 51 (4):412–413, April 2004. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Slade:2002:SLS

[Sla02] Gordon Slade. Scaling limits and super-Brownian motion. Notices of the American Mathematical Society, 49(9):1056– 1067, October 2002. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sloane:2003:LEI

[Slo03] N. J. A. Sloane. The on-line encyclopedia of integer sequences.

Notices of the American Mathematical Society, 50(8):912–915, September 2003. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Springer:2000:HDK

[Spr00] Tonny A. Springer. H. D. Kloosterman and his work. Notices of the American Mathematical Society, 47(8):862–867, September 2000. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Scott:2005:WF

[SST^{+}05] Leonard Scott, Ronald Solomon, John Thompson, John Walter, and Efim Zelmanov. Walter Feit (1930–2004). Notices of the American Mathematical Society, 52(7):728–735, August 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Stasheff:2004:WLO

[Sta04] Jim Stasheff. What is … an operad? Notices of the American Mathematical Society, 51 (6):630–631, June/July 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Stephenson:2003:CPM

[Ste03] Kenneth Stephenson. Circle packing: a mathematical tale. Notices of the American Mathematical Society, 50(11):1376– 1388, December 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Steel:2007:WLW

[Ste07] John R. Steel. What is … a Woodin cardinal? Notices of the American Mathematical Society, 54(9):1146–1147, October 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Stillwell:2001:SC

[Sti01] John Stillwell. The story of the 120-cell. Notices of the American Mathematical Society, 48 (1):17–24, January 2001. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Stopple:2006:NDH

[Sto06]  Jeffrey Stopple.  Notes on

the Deuring–Heilbronn phenomenon. Notices of the American Mathematical Society, 53 (8):864–875, September 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sturmfels:2005:WLG

[Stu05] Bernd Sturmfels. What is … a Gröbner basis? Notices of the American Mathematical Society, 52(10):1199–1200, November 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sunada:2008:CNM

[Sun08] Toshikazu Sunada. Crystals that nature might miss creating. Notices of the American Mathematical Society, 55(2):208–215, February 2008. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sussmann:2009:BAP

[Sus09] Hector Sussmann. The best of all possible worlds: mathematics and destiny [book review of MR2259005]. Notices of the American Mathematical Society, 56(3):376–379, March 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Szpiro:2003:KCH

[Szp03] George Szpiro, editor. Kepler’s Conjecture: how some of the greatest minds in history helped solve one of the oldest math problems in the world. Wiley, New York, NY, USA, 2003. ISBN 0-471-08601-0. viii + 296 pp. LCCN QA93 .S97

2003. URL http://www.loc.
gov/catdir/bios/wiley046/
2002014422.html;   http://
www.loc.gov/catdir/description/
wiley039/2002014422.html;
 http://www.loc.gov/catdir/
toc/wiley031/2002014422.html.

Szpiro:2007:ISS

[Szp07] George Szpiro. Interview with Stephen Smale. Notices of the American Mathematical Society, 54(8):995–997, September 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tanton:2006:MCO

[Tan06] James Tanton. Math circles and olympiads. MSRI asks: is the U.S. coming of age? Notices of the American Mathematical Society, 53(2):200–205, February 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tao:2001:RNS

[Tao01] Terence Tao. From rotating needles to stability of waves: emerging connections between combinatorics, analysis, and PDE. Notices of the American Mathematical Society, 48 (3):294–303, March 2001. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Toledo:2008:JE

[Tol08] Domingo Toledo. James Eells 1926–2007. Notices of the American Mathematical Society, 55 (6):704–706, June/July 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Treves:2003:LS

[TPY03] François Treves, Gilles Pisier, and Marc Yor. Laurent Schwartz (1915–2002). Notices of the American Mathematical Society, 50(9):1072–1084, October 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Treves:2000:TTG

[Tre00] François Treves. A treasure trove of geometry and analysis: the hyperquadric. Notices of the American Mathematical Society, 47(10):1246– 1256, November 2000. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Terng:2000:GS

[TU00] Chuu-Lian Terng and Karen Uhlenbeck. Geometry of solitons. Notices of the American Mathematical Society, 47(1):17– 25, January 2000. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Tu:2006:LWR

[Tu06] Loring W. Tu. The life and works of Raoul Bott. Notices of the American Mathematical Society, 53(5):554–570, May 2006. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Vakil:2003:MSC

[Vak03] Ravi Vakil. The moduli space of curves and its tautological ring. Notices of the American Mathematical Society, 50(6):647–658, June/July 2003. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Vershik:2007:MWF

[VBKR07] A. M. Vershik, Jean Bourgain, Harry Kesten, and Nicolai Reshetikhin. The mathematical work of the 2006 Fields medalists. Notices of the American Mathematical Society, 54 (3):388–404, March 2007. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Vogan:2007:CT

[Vog07] David Vogan. The character table for E_{8}. Notices of the American Mathematical Society, 54 (9):1122–1134, October 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Vogtmann:2008:WLO

[Vog08] Karen Vogtmann. What is … outer space? Notices of the American Mathematical Society, 55(7):784–786, August 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Willinger:2009:MIS

[WAD09] Walter Willinger, David Alderson, and John C. Doyle. Mathematics and the Internet: a source of enormous confusion and great potential. Notices of the American Mathematical Society, 56(5):586–599, May 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Washington:2008:EHF

[Was08] Talitha M. Washington. Evansville honors the first black Ph.D. in mathematics and his family. Notices of the American Mathematical Society, 55(5):588–589, May 2008. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Weller:2004:II

[WBD^{+}04] Kirk Weller, Anne Brown, Ed Dubinsky, Michael McDonald, and Cynthia Stenger. Intimations of infinity. Notices of the American Mathematical Society, 51(7):741–750, August 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Weeks:2004:PDS

[Wee04] Jeffrey Weeks. The Poincaré dodecahedral space and the mystery of the missing fluctuations. Notices of the American Mathematical Society, 51(6):610–619, June/July 2004. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wells:2007:ATF

[Wel07] Benjamin Wells. Alfred Tarski, friend and daemon. Notices of the American Mathematical Society, 54(8):982–984, September 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wiegand:2006:WLS

[Wie06] Roger Wiegand. What is … a syzygy? Notices of the American Mathematical Society, 53 (4):456–457, April 2006. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wiedijk:2008:FPG

[Wie08] Freek Wiedijk. Formal proof — getting started. Notices of the American Mathematical Society, 55(11):1408–1417, December 2008. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Wigderson:2003:WMS

[Wig03] Avi Wigderson. On the work of Madhu Sudan. Notices of the American Mathematical Society, 50(1):45–50, January 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Woodin:2001:CH

[Woo01a] W. Hugh Woodin. The continuum hypothesis. I. Notices of the American Mathematical Society, 48(6):567–576, June/ July 2001. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Woodin:2001:CHI

[Woo01b] W. Hugh Woodin. The continuum hypothesis. II. Notices of the American Mathematical Society, 48(7):681–690, August 2001. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). See correction [Woo02].

Woodin:2002:CCH

[Woo02] W. Hugh Woodin. Correction:

“The continuum hypothesis. II”. Notices of the American Mathematical Society, 49(1):46, January 2002. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic). See [Woo01b].

Wright:2004:DCG

[Wri04] David J. Wright. Double cusp group. Notices of the American Mathematical Society, 51 (11):1332–1333, December 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Yong:2007:WLY

[Yon07] Alexander Yong. What is … a Young tableau? Notices of the American Mathematical Society, 54(2):240–241, February 2007. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zelevinsky:2007:WLC

[Zel07] Andrei Zelevinsky. What is … a cluster algebra? Notices of the American Mathematical Society, 54(11):1494–1495, December 2007. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ziegler:2004:GPN

[Zie04] Günter M. Ziegler. The great prime number record races. Notices of the American Mathematical Society, 51(4):414–416, April 2004. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zitarelli:2009:CSA

[Zit09] David E. Zitarelli. Connected sets and the AMS, 1901–1921.

Notices of the American Mathematical Society, 56(4):450–458, April 2009. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).